### v1

# A new expansion of the Riemann zeta function

Abstract
After a brief introduction to Ramanujan’s method of summation, we give an expansion of the Riemann Zeta function in the critical strip as a convergent series $\sum_{m\geq 0}x_{m}P_{m}(s)$ where the functions $P_{m}$ are polynomials with their roots on the line $\{\Re(s)=1/2\}$, the coefficients $x_{m}$ being finite linear combinations of the Euler constant $\gamma$ and the values $\zeta(2),\zeta(3),\dots,\zeta(m+1).$
1 Introduction
There are many polynomial expansions of the Riemann zeta function. Some of them (cf.[E-K-K], [F-V],[B],[M], [V]) are based on the Newton interpolation formula for functions related to Zeta. For example
$\zeta(s)-\frac{1}{s-1}=\sum_{m=0}^{+\infty}(-1)^{m}c_{m}\binom{s-1}{m}$,
with
$c_{m}=\gamma+\sum_{k=1}^{m}\binom{m}{k}(-1)^{k}\big(\zeta(k+1)-\frac{1}{k}\big).$
Others are linked to the well known expansion
$\zeta(s)-\frac{1}{s-1}=\sum_{m=0}^{+\infty}(-1)^{m}\gamma_{m}\frac{(s-1)^{m}}{m!}$
where $\gamma_{m}$ are the Stieltjes constants. This gives (cf.[E-G])
$\zeta(s)-\frac{s}{s-1}=\sum_{m=0}^{+\infty}(-1)^{m}l_{m}(\frac{s-1}{s})^{m},$
with
$l_{m}=\sum_{k=1}^{m}\binom{m-1}{k-1}(-1)^{m-k}\frac{\gamma_{k}}{k!}.$
and for $s=\frac{1}{2}+it$ $(t\in\mathbb{R}$) one obtain an expansion of $\zeta(\frac{1}{2}+it)$ on the orthonormal basis $(e_{m})$ of $L^{2}(\mathbb{R},\frac{dt}{2\pi(t^{2}+1/4)})$ given by
$e_{m}(t)=\Big(\frac{1-2it}{1+2it}\Big)^{m}.$
There is also (cf. [K]) an expansion of the function
$\Xi(t)=\frac{s(s-1)}{2}\pi^{-s/2}\Gamma(s/2)\zeta(s),\ \ \ s=\frac{1}{2}+it,\ (t\in\mathbb{R}),$
which takes the form
$\Xi(t)=e^{\pi t/4}\sum_{m=0}^{+\infty}b_{m}p_{m}(t),$
where $p_{m}(t)$ are polynomials related to the Mellin transform of odd Hermite functions. These polynomials are orthogonal in $(L^{2}(\mathbb{R},|\Gamma(3/4+it/2)|^{2}dt)$, but it seems that the coefficients $b_{m}$ does not have a simple form.
The polynomial expansion we give in this paper is related to the Mellin transform of the functions $x\mapsto e^{-x}L_{m}(2x)$ where $(L_{m})_{m\geq 0}$ are the Laguerre polynomials defined by
$L_{m}(x)=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}\frac{x^{k}}{k!}.$
The generating function of the Laguerre polynomials (cf. [Leb]) gives for $x\in[0,+\infty[$
$e^{-xz}=\sum_{m=0}^{+\infty}2e^{-x}L_{m}(2x)\psi_{m}(z).$
where $(\psi_{m})_{m\geq 0}$ the sequence of the functions defined in the half-plane $\{Re(z)>0\}$ by
$\psi_{m}(z)=\frac{1}{1+z}\big(\frac{z-1}{z+1}\big)^{m}.$
By Mellin transform of the function $x\mapsto e^{-xz}$, we obtain for $\Re(z)>0$ and $\Re(s)>0$
$\Gamma(s)z^{-s}=\sum_{m=0}^{+\infty}\Big(2\int_{0}^{+\infty}e^{-x}L_{m}(2x)x^{s-1}\ dx\Big)\psi_{m}(z),$
Since we can write
$2\int_{0}^{+\infty}e^{-x}L_{m}(2x)x^{s-1}\ dx={\Gamma(s)}P_{m}(s),$
where $P_{m}$ is the polynomial
$P_{m}(s)=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k+1}\ \frac{\Gamma(s+k)}{\Gamma(s)}\frac{1}{k!},$
then we have, for $\Re(s)>0$, the expansion
$z^{-s}=\sum_{m=0}^{+\infty}\psi_{m}(z)P_{m}(s).$
Take $z=1,2,3,\dots$, then we get an infinite sequence of relations that it’s tempting to sum up, to obtain
$\sum_{n\geq 1}\frac{1}{n^{s}}=\sum_{m=0}^{+\infty}\Big(\sum_{n\geq 1}\psi_{m}(n)\Big)P_{m}(s).$
In section 3, we prove that, for $0<\Re(s)<1$, this last formula is correct if we replace the divergent series $\sum_{n\geq 1}$ by their Ramanujan’s summations $\sum_{n\geq 1}^{\mathcal{R}}$, and then we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\sum_{m=0}^{+\infty}z_{m}P_{m}(s)\ \text{ where }\ z_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}.$
In section 2 we give a short introduction to Ramanujan’s summation of divergent series. Since we will see that, for $s\neq 1$, we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\zeta(s)-\frac{1}{s-1},$
then, after an explicit evaluation of $z_{n}$, we obtain, for $0<\Re(s)<1$, the following expansion
$\zeta(s)=\sum_{m=0}^{+\infty}x_{m}P_{m}(s)$
with
$x_{m}=\gamma+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\zeta(k+1)+r_{m}$
and $r_{m}$ are rational numbers explicitely given. For $s=\frac{1}{2}+it$ with $t\in\mathbb{R}$, the preceding relation gives the expansion
$\zeta(\frac{1}{2}+it)=2\sum_{m=0}^{+\infty}x_{m}Q_{m}(t)$
where the polynomials
$Q_{m}:t\mapsto\frac{1}{2}P_{m}(\frac{1}{2}+it)$
form complete orthonormal system in $\big(L^{2}(-\infty,+\infty),\frac{dt}{\cosh(\pi t)}\big)$. Thus, for every $m\geq 1$, the roots of $Q_{m}$ are real, and this implies that the roots of $P_{m}$ are on the line $\{\Re(s)=\frac{1}{2}\}$.
2 Ramanujan’s summation
2.1 Ramanujan’s constant of a series
Let a function $f\in C^{1}(]0,+\infty[)$, the first order Euler-MacLaurin formula says that, for every integer $n\geq 1$
$\sum_{k=1}^{n}f(k)=\int_{1}^{n}f(x)\,dx+\frac{1}{2}\big(f(n)+f(1)\big)+\int_{1}^{n}\big(x-[x]-\frac{1}{2}\big)f^{\prime}(x)\ dx$}, (1)
where $[x]$ denote the integer part of $x$. For $x\geq 1$, let
$b(x)=x-[x]-\frac{1}{2}$
and let’s assume that the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\ dx$ is convergent. The previous formula gives
$\sum_{k=1}^{n}f(k)=\int_{1}^{n}f(x)\,dx+C(f)+\frac{1}{2}f(n)-\int_{n}^{+\infty}b(x)f^{\prime}(x)\ dx$}, (2)
 $\text{ with }\ \ C(f)=\frac{1}{2}f(1)+\int_{1}^{+\infty}b(x)f^{\prime}(x)\ dx.$
The constant $C(f)$ is called Ramanujan’s constant of the series $\sum_{n\geq 1}f(n)$ (cf. [H1]).
Clearly $C(f)$ depends linearly of $f$. And if the series $\sum_{n\geq 1}f(n)$ is convergent, then by (2) we have
$C(f)=\sum_{k=1}^{+\infty}f(k)-\int_{1}^{+\infty}f(x)\ dx.$
We observe that $C(f)$ is also well defined for some divergent series. We have immediately $C(1)=1/2.$ And applying (2) to the function $f:x\mapsto 1/x$, we have for every integer $n\geq 1$
$\sum_{k=1}^{n}\frac{1}{k}=\log n+C(f)+\frac{1}{2n}+\int_{n}^{+\infty}b(x)\frac{1}{x^{2}}\ dx,$
which gives that the Ramanujan’s constant of the series $\sum_{n\geq 1}\frac{1}{n}$ is the Euler constant :
$C(f)=\gamma=\lim_{n\to+\infty}\Big(\sum_{k=1}^{n}\frac{1}{k}-\log n\Big).$
2.2 Plana’s formula
The following summation formula was discovered by Plana in 1820 (and also given by Abel in 1823) (cf. [Lin]).
Theorem 1.
Let $f$ be an analytic function in the half-plane $\{\Re(z)>0\}$ such that there is a constant ${a}<2\pi$ and, for any interval $c,d]\subset]0,+\infty[$, a constant $C_{c,d}>0$ with
$\left|f(x+it)\right|\leq C_{c,d}\ e^{a|t|},\ \text{ for }x\in[c,d]\ \text{ and all }t\in\mathbb{R}.$
Then we have Plana’s formula : for any integer $n\geq 1$,
$\sum_{k=1}^{n}f(k)$ $=\frac{f(1)+f(n)}{2}+\!\int_{1}^{n}f(u)\ du$
$+i\!\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\,\ dt-i\!\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\,\ dt\,.$ (3)
If we now compare Euler-MacLaurin formula (1) and Plana’s (3), then we get, for $n\geq 1$,
$\int_{1}^{n}b(x){f^{\prime}(x)}\ dx=i\!\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\ dt-i\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\,\ dt\,.$
If the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\,dx$ is convergent and if
$\lim_{n\to+\infty}\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\ dt=0,$
then we have
$\int_{1}^{+\infty}b(t){f^{\prime}(t)}\ dt=i\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\,\ dt.$
Thus, in this case the Ramanujan’s constant $C(f)$, previously defined by (2), is also
$C(f)=\frac{1}{2}f(1)+i\int_{0}^{+\infty}\frac{{f(1+it)-f(1-it)}}{e^{2\pi t}-1}\,\ dt.$ (4)
Take now this last formula as the definition of Ramanujan’s constant $C(f)$ for a function $f$ verifying the hypotheses of theorem 1.
Remark 1.
Let $f$ and $g$ two functions verifying the assumptions of theorem 1 and such that $f(n)=g(n)$ for every integer $n\geq 1$. Then Ramanujan’s constants $C(f)$ and $C(g)$, defined by (4), need not necessarily be the same. For example, if $f\equiv 0$ and $g:z\mapsto\sin(\pi z)$, then $C(f)=0$, but
$C(g)=i\int_{0}^{+\infty}\frac{\sin\big(\pi(1+it)\big)-\sin\big(\pi(1-it)\big)}{e^{2\pi t}-1}\,\ dt=\int_{0}^{+\infty}\frac{e^{\pi t}-e^{-\pi t}}{e^{2\pi t}-1}\,\ dt\;\;=\;\int_{0}^{+\infty}e^{-\pi t}\ dt=\frac{1}{\pi}.$
To avoid functions like $z\mapsto\sin(\pi z)$, we will consider a space of analytic functions such that Ramanujan’s constant, defined by (4), depends only on the values $f(n)$ for all integers $n\geq 1$}.
2.3 The space $\mathcal{E}^{\pi}$
Definition 1.
Let ${\alpha}>0$ and ${\mathcal{E}}^{\alpha}$ the space of analytic functions $f$ in the half plane $\{\Re(z)>0\}$ such that there is some constants $\emph{a}<\alpha$ and $\emph{b}>0$ with :
for every ${\varepsilon}>0$, there is a constant $C_{\varepsilon}$ such that
$\left|f(z)\right|\;\;\leq\;\;C_{\varepsilon}\ e^{b\,\Re(z)}\ e^{a\,|\Im(z)|},\ \text{ if }\ \Re(z)>{\varepsilon}.$ (5)
Example 1.
Let $f_{s}\;:\;z\mapsto e^{sz}$, with $s=\sigma+i\tau$, $\sigma\in\mathbb{R}$ and $\tau\in\mathbb{R}$ . For $\Re(z)>0$, we have
$|f_{s}(z)|\;\;=\,\;e^{\sigma\Re(z)-\tau\Im(z)}.$
Thus $f_{s}\in\mathcal{E}^{\alpha}$ if and only if $\alpha<\Im(s)<\alpha$.
Properties
a) If $f\in\mathcal{E}^{\alpha}$ and $g\in\mathcal{E}^{\beta}$, alors $fg\in E^{\alpha+\beta}$.
b) If $f\in\mathcal{E}^{\alpha}$ and if $F(z)=\int_{1}^{z}f(u)du$ then $F\in\mathcal{E}^{\alpha}$.
An important theorem for the rest of this section is the following.
Theorem 2.
Carlson’s theorem (cf. [H2])
If $f\in\mathcal{E}^{\pi}$ is such that $f(n)=0$ for all $n\;=\;1,\,2,\,\ldots$, then $f\equiv 0$.
Thus a function $f\in\mathcal{E}^{\pi}$ is only determined by the values $f(1),f(2),f(3),\dots$.
Remark 2.
The function $g\;:\;z\mapsto\sin(\pi z)$ is such that $g(n)=0$ for every integer $n\geq 1$}, but $g\notin\mathcal{E}^{\pi}$ since for $x\in\mathbb{R}$ and $t\in\mathbb{R}$
$|g(x+it)|\sim\frac{1}{2}e^{\pi|t|}\text{ when }t\to\infty$
2.4 Ramanujan’s method of summation
Let $f$ and $g$ two functions verifying the hypothesis of theorem 1. By Carlson’s theorem we know that if $f\in\mathcal{E}^{\pi}$ and $g\in\mathcal{E}^{\pi}$ verify f(n)=g(n){f(n)=g(n)} for $n=1,2,3,\dots$, then for x>0{x>0} and $t\in\mathbb{R}$, we have
f(x+it)f(xit)=g(x+it)g(xit),{f(x+it)-f(x-it)}={g(x+it)-g(x-it)},
thus, by (4), we have $C(f)=C(g).$ This leads to the following definition.
Definition 2.
For $f\in\mathcal{E}^{\pi}$, we define the Ramanujan’s summation of the series $\sum_{n\geq 1}f(n)$ by
$\sum_{n\geq 1}^{\mathcal{R}}f(n)=\frac{f(1)}{2}+i\!\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\,\ dt\,.$ (6)
It is clear that $\sum_{n\geq 1}^{\mathcal{R}}f(n)$ depends linearly on $f$.
Remark 3.
Let $f\in\mathcal{E}^{\pi}$ then by (3), we have for every integer $n\geq 1$
$\sum_{k=1}^{n}f(k)=\!\int_{1}^{n}f(u)\ du+\frac{f(n)}{2}+\sum_{n\geq 1}^{\mathcal{R}}f(n)-i\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\,\ dt\,.$ (7)
If $\lim_{n\to+\infty}f(n+it)=0$, for every t\ $\in\mathbb{R}, t$hen, by dominated convergence, we get
$\lim_{n\to+\infty}\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\,\ dt\,=0,$
and if in addition, the integral $\int_{1}^{+\infty}f(u)\ du$ is convergent then the series $\sum_{n\geq 1}f(n)$ is convergent. In this case we have by (7)
$\sum_{n\geq 1}^{\mathcal{R}}f(n)\;\;=\sum_{n=1}^{+\infty}f(n)-\!\!\int_{1}^{+\infty}\!\!f(x)\ dx.$
Example 2.
For $f:x\mapsto\frac{1}{x^{s}}$, with $\Re(s)>1$, we get
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}\;\;=\sum_{n=1}^{+\infty}\frac{1}{n^{s}}-\!\!\int_{1}^{+\infty}\!\!\frac{1}{x^{s}}\ dx,$
thus, for $\Re(s)>1,$ we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}\;\;=\zeta(s)-\frac{1}{s-1}.$ (8)
2.5 Interpolation of partial sums
To evaluate Ramanujan’s summation of a divergent series $\sum_{n\geq 1}f(n)$, a simpler way than the formula (6) can be obtained using an interpolation function of partial sums of this series . In chapter VI of his Notebooks, Ramanujan uses, without any definition, such a function $\varphi$ which satisfies
${\varphi(x)-\varphi(x-1)=f(x)}\ \text{ with }\varphi(0)=0.$
The following theorem clarifies this point.
Theorem 3.
Let $f\in\mathcal{E}^{\pi}$. There is a unique function $\varphi_{f}\in\mathcal{E}^{\pi}$, such that
[…]
Proof.
a) Existence. Since $f\in\mathcal{E}^{\pi}$ then the function $\varphi_{f}$, defined for $\Re(z)>0$, by
$\varphi_{f}(z)=\int_{1}^{z}f(u)\ du+\frac{f(z)}{2}+\sum_{n\geq 1}^{\mathcal{R}}f(n)-i\int_{0}^{+\infty}\frac{f(z+it)-f(z-it)}{e^{2\pi t}-1}\ dt,$ (10)
is also in $\mathcal{E}^{\pi}$. By (3) we have $\varphi_{f}(n)=\sum_{k=1}^{n}f(k)$ for every integer $n\geq 1$, thus the function $\varphi_{f}$ verifies
$\varphi_{f}(n+1)-\varphi_{f}(n)\,=f(n+1),\text{ for }n=1,2,3,\dots.$
Since the functions $z\mapsto\varphi_{f}(z+1)-\varphi_{f}(z)$ and $z\mapsto f(z+1)$ are in $\mathcal{E}^{\pi}$, then by Carlson’s theorem, we get
$\varphi_{f}(z+1)-\varphi_{f}(z)=f(z+1),\quad\!\!\text{ for all }z\text{ such that }\Re(z)>0.\$(11)
The condition $\varphi_{f}(1)=f(1)$ is simply a consequence of (6) and (10).
b) Uniqueness. If $\varphi\in\mathcal{E}^{\pi}$ is such that
$\varphi(x+1)-\varphi(x)=f(x+1)\ \text{ for }x>0,\ \ \text{ with }\varphi(1)=f(1),$
then, for every integer $n\geq 1$, we get
$\varphi(n)-f(1)=\varphi(n)-\varphi(1)\;=\sum_{k=1}^{n-1}\big(\varphi(k+1)-\varphi(k)\big)=\sum_{k=1}^{n-1}f(k+1),$
thus $\varphi(n)=\sum_{k=1}^{n}f(k)=\varphi_{f}(n)$ for every integer $n\geq 1$. Since $\varphi$ and $\varphi_{f}$ are in the space $\mathcal{E}^{\pi}$, then by Carlson’s theorem we get $\varphi=\varphi_{f}.$
Remark 4.
The preceding proof asserts that if $f\in\mathcal{E}^{\pi}$ then
$\varphi_{f}(n)=\sum_{k=1}^{n}f(k)\text{ for every integer }n\geq 1.$
The function $\varphi_{f}$ given by (10) is the unique function in $\mathcal{E}^{\pi}$ wich interpolates the partial sums of the series $\sum_{n\geq 1}f(n)$. Note that, by (11), we have for $\Re(z)>0$
$\varphi_{f}(z)=\varphi_{f}(z+1)-f(z+1),$
and since the right-hand side defines an analytic function in the half-plane $\{\Re(z+1)>0\}$, this relation gives an analytic continuation of $\varphi_{f}$ in the half-plane $\{\Re(z)>-1\}.$ Let’s denote again by $\varphi_{f}$ this continuation, then the relation $\varphi_{f}(1)=f(1)$} provides the condition required by Ramanujan
$\varphi_{f}(0)=\varphi_{f}(1)-f(1)=0.$}
Example 3.
Let $f:z\mapsto 1/z$, then $\varphi_{f}$ is related to digamma function which is defined, for $\Re(z)>0$, by
$\Psi(z)=\frac{\Gamma^{\prime}(z)}{\Gamma(z)}.$
This function is analytic in $\{\Re(z)>0\}$, and by logarithmic differentiation of the classical product formula for $\Gamma$ (cf. [G-R]) we get for $\Re(z)>0$
$\Psi(z)=-\frac{1}{z}-\gamma+\sum_{n=1}^{+\infty}\big(\frac{1}{n}-\frac{1}{n+z}\big)=-\frac{1}{z}-\gamma+\sum_{n=1}^{+\infty}\frac{z}{n(n+z)}$ (12)
Then it is easily verified that $\Psi\in\mathcal{E}^{\pi}$ and that
$\Psi(z+2)\;-\;\Psi(z+1)=\frac{1}{z+1}$
Since $\Psi(1)=-\gamma$, then, by the previous theorem, we get for $\Re(z)>0$
$\varphi_{f}(z)=\Psi(z+1)+\gamma.$
A very useful formula for Ramanujan’s summation is provided by the following theorem.
Theorem 4.
Let $f\in\mathcal{E}^{\pi}$, and let $\varphi\in\mathcal{E}^{\pi}$ be a solution of
$\varphi(x+1)-\varphi(x)$ $=$ $f(x+1),\ \text{ for all }\ x>0,$
$\varphi(1)$ $=$ $f(1).$
Then we have
$\sum_{n\geq 1}^{\mathcal{R}}f(n)\;\;=\int_{0}^{1}\varphi(x)dx.$
Proof.
By theorem 3, we have $\varphi=\varphi_{f}$. Let $F(z)=\int_{1}^{z}f(u)\ du$, then integration of (10) gives
$\int_{1}^{2}\varphi_{f}(x)\ dx\;\;=\int_{1}^{2}F(u)\ du\,+\,\frac{1}{2}\int_{1}^{2}f(x)\ dx\,+\sum_{n\geq 1}^{\mathcal{R}}f(n)-iD(f),$ (13)
where
$D(f)=\int_{1}^{2}\Big(\int_{0}^{+\infty}\frac{f(x+it)-f(x-it)}{e^{2\pi t}-1}\,\ dt\Big)\ dx.$
By Fubini’s theorem we have
$D(f)\;\;=\int_{0}^{+\infty}{\Big(\int_{1}^{2}{f(x+it)-f(x-it)}\ dx\Big)}\frac{1}{e^{2\pi t}-1}dt,$
thus
$D(f)=\int_{0}^{+\infty}\frac{F(2+it)-F(2-it)}{e^{2\pi t}-1}dt-\!\int_{0}^{+\infty}\frac{F(1+it)-F(1-it)}{e^{2\pi t}-1}dt.$
If we apply Plana’s formula to the function $F$, with $n=2$, then we get
$-\big(\frac{1}{2}F(1)+\frac{1}{2}F(2)\big)+\!\int_{1}^{2}F(t)dt$ $=i\int_{0}^{+\infty}\frac{F(2+it)-F(2-it)}{e^{2\pi t}-1}dt$
$-i\int_{0}^{+\infty}\frac{F(1+it)-F(1-it)}{e^{2\pi t}-1}dt,$
which is
$iD(f)=-\frac{F(2)}{2}+\int_{1}^{2}F(t)\ dt.$
Then (13) becomes
$\int_{1}^{2}\varphi_{f}(x)\ dx$ $=$ $\frac{1}{2}\int_{1}^{2}f(x)\ dx+\sum_{n\geq 1}^{\mathcal{R}}f(n)+\frac{F(2)}{2}$
$=$ $\int_{1}^{2}f(x)\ dx+\sum_{n\geq 1}^{\mathcal{R}}f(n).$
from which we deduce that
$\sum_{n\geq 1}^{\mathcal{R}}f(n)=\int_{0}^{1}\big(\varphi_{f}(x+1)-f(x+1)\big)\ dx=\int_{0}^{1}\varphi_{f}(x)\ dx.$
Corollary 1.
Let $f\in\mathcal{E}^{\pi}$, then the function $x\mapsto f(x+1)$ belongs to $\mathcal{E}^{\pi}$ and we have the shift formula
$\sum_{n\geq 1}^{\mathcal{R}}f(n+1)\;\;=\sum_{n\geq 1}^{\mathcal{R}}f(n)-f(1)+\int_{0}^{1}f(x+1)\ dx.$
Proof.
Let $g:x\mapsto f(x+1)$, then it is easily verified, by theorem 3, that we have for $x>0$
$\varphi_{g}(x)=\varphi_{f}(x+1)-f(1).$
Thus we have
$\int_{0}^{1}\varphi_{g}(x)dx=\int_{0}^{1}\varphi_{f}(x+1)dx-f(1)=\int_{0}^{1}\varphi_{f}(x)dx+\int_{0}^{1}f(x+1)dx-f(1),$
and the conclusion follows from theorem 4. ∎
Example 4.
Let $f:x\mapsto 1/x$. By example 3, we have $\varphi_{f}(x)=\Psi(x+1)+\gamma,$ thus we get
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}=\int_{0}^{1}(\Psi(x+1)+\gamma)dx=\int_{0}^{1}\frac{\Gamma^{\prime}}{\Gamma}(x+1)dx+\gamma,$
thus we verify again that
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}=\gamma,$ (14)
Observe that, by corollary 1, we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}=\gamma-1+\log(2).$
Example 5.
Since, fro any $x>0$, we have
$\log(\Gamma(x+2))-\log(\Gamma(x+1))=\log(x+1)$
and $\log(\Gamma(2))=\log(1)$, we deduce that $\varphi_{\log}(x)=\log(\Gamma(x+1))$. Now if we use the Raabe’s formula, that is :
$\int_{a}^{a+1}\log(\Gamma(x))dx=a\log(a)-a+\log(\sqrt{2\pi}),$
then we get
$\sum_{n\geq 1}^{\mathcal{R}}\log(n)=\log(\sqrt{2\pi})-1.$
Example 6.
Let $f(x)=\Psi(x)$ and $\varphi(x)=x\Psi(x)-x+1$. We have $f(1)=\Psi(1)=\varphi(1)$ and
$\varphi(x+1)-\varphi(x)=\Psi(x+1)=f(x+1).$
Thus, by the previous theorem, we get
$\sum_{n\geq 1}^{\mathcal{R}}\Psi(n)=\int_{0}^{1}x\Psi(x)dx+\frac{1}{2}=-\int_{0}^{1}\log(\Gamma(x))dx+\frac{1}{2},$
then by Raabe’s formula
$\sum_{n\geq 1}^{\mathcal{R}}\Psi(n)=-\log(\sqrt{2\pi})+\frac{1}{2}.$
Example 7.
We can evaluate $\sum_{n\geq 1}^{\mathcal{R}}n^{k}$ for $k=1,2,3,\dots$ by means of Bernoulli numbers. Let the Bernoulli polynomials $B_{n}(x)$ be defined by $B_{0}(x)=1$ and, for $k\geq 1,$
$B^{\prime}_{k}(x)=nB_{k-1}(x)\ \text{ with }\ \int_{0}^{1}B_{k}(x)dx=0.$
The Bernoulli numbers are the numbers $B_{k}=B_{k}(0)$. These polynomials verify (cf. [G-R])
$B_{k+1}\big((x+1)+1\big)-B_{k+1}(x+1)=(k+1)(x+1)^{k}.$
For $k\geq 1$, we have $B_{k+1}(1)=B_{k+1}$. Let the function $f:x\mapsto x^{k}$ with an integer $k\geq 1$, then by theorem 3, we obtain
$\varphi_{f}(x)=\frac{B_{k+1}(x+1)}{k+1}-\frac{B_{k+1}}{k+1}.$
Since
$\int_{0}^{1}\frac{B_{k+1}(x+1)}{k+1}dx\;\;=\int_{0}^{1}\frac{B_{k+1}(x)}{k+1}dx\,+\int_{0}^{1}x^{k}dx=\frac{1}{k+1},$
then theorem 4 gives
$\sum_{n\geq 1}^{\mathcal{R}}n^{k}=\frac{1-B_{k+1}}{k+1}.$ (15)
Remark 5.
Relations (14) and (15) can be written in terms of the Zeta function :
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}$ $=$ $\lim_{s\to 1}\Big(\zeta(s)-\frac{1}{s-1}\Big),$
$\sum_{n\geq 1}^{\mathcal{R}}n^{k}$ $=$ $\zeta(-k)-\frac{1}{-k-1}\text{ for every integer }k\geq 1.$
Note also that
$\sum_{n\geq 1}^{\mathcal{R}}n^{0}=\frac{1}{2}=\zeta(0)+1,$
and by (9) we have in the half-plane $\{\Re(s)>1\}$
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\zeta(s)-\frac{1}{s-1}.$ (16)
All these formulae are special cases the following general result.
Theorem 5.
The function
$s\mapsto\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}\;\;$
is analytic on $\mathbb{C}$ and we have for every $s\in\mathbb{C}\setminus\{1\}$
$\zeta(s)=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}+\frac{1}{s-1}$ (17)
Proof.
First, it is easily seen that, for every $s\in\mathbb{C}$, the function $z\mapsto 1/{z^{s}}$ is in the space $\mathcal{E}^{\alpha}$, for all $\alpha>0.$ Since, if $s=\sigma+i\tau$ ($\sigma\in\mathbb{R},\ \tau\in\mathbb{R}$), we can write
$\Big|\frac{1}{z^{s}}\Big|=\Big|\frac{1}{(x+it)^{s}}\Big|=\Big|e^{-s\log(x+it)}\Big|=e^{-{\sigma}\log\sqrt{x^{2}+t^{2}}+\tau\arctan(\frac{t}{x})}\leq e^{|\tau|\pi/2}(x^{2}+t^{2})^{-\sigma/2}$
where $z=x+it$, with $x>\varepsilon>0$ and $t\in\mathbb{R}$. Thus the sum $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}$ is well defined and, by (6), we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\;\;\frac{1}{2}+i\int_{0}^{+\infty}\Big({\frac{1}{(1+it)^{s}}-\frac{1}{(1-it)^{s}}}\big)\frac{1}{e^{2\pi t}-1}\,\ dt\,$
$\frac{1}{2}+2\int_{0}^{+\infty}\frac{\sin(s\arctan(t))}{(1+t^{2})^{s/2}}\frac{1}{e^{2\pi t}-1}dt.$
This last integral depends analyticaly of $s\in\mathbb{C}$ so the function $s\mapsto\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}$ is analytic on $\mathbb{C}$. Since the function $s\mapsto\zeta(s)-\frac{1}{s-1}$ is analytic on $\mathbb{C}\setminus\{1\}$ then, by analytic continuation of (16), we get
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\zeta(s)-\frac{1}{s-1}\text{ for any }s\in\mathbb{C}\setminus\{1\}.$
3 An expansion of Zeta in the critical strip
We are now going to use expression (17) to obtain, in the strip $\{0<\Re(s)<1\}$, a polynomial expansion of Zeta
$\zeta(s)=\sum_{m=0}^{+\infty}x_{m}P_{m}(s).$
As we have seen in the Introduction the polynomials $P_{m}$ are related to the Laguerre polynomials by Mellin transform. We now give another approach to these polynomials.
3.1 The sequence of polynomials $(P_{m})$
We start with the observation that for $s\in\mathbb{C}$, and for every $n>0$, we have
$\frac{1}{n^{s}}=\frac{1}{n+1}G_{s}(\frac{n-1}{n+1}),$ (18)
where $G_{s}$ is the function defined on $D(0,1)=\{z\in\mathbb{C};|z|<1\}$ by
$G_{s}:z\mapsto\frac{2}{1-z}\Big({\frac{1-z}{1+z}}\Big)^{s}$ (19)
Since $\Re\big({\frac{1-z}{1+z}}\big)>0$ for $z\in D(0,1)$, we see that the function $G_{s}$ is analytic on $D(0,1)$. Let $P_{m}(s)$ denote the coefficients of its Taylor expansion
$\frac{2}{1-z}\Big({\frac{1-z}{1+z}}\Big)^{s}=\sum_{m=0}^{+\infty}P_{m}(s)z^{m}\ \text{ for }\ \ \ z\in D(0,1)$ (20)
Let $z=\frac{x}{1+x}$, with $x\in]0,1[$, then, using (20), we get
$\sum_{m=0}^{+\infty}P_{m}(s)\Big(\frac{x}{1+x}\Big)^{m+1}=2x(1+2x)^{-s}=\sum_{k=0}^{+\infty}(-1)^{k}2^{k+1}\frac{(s)_{k}}{k!}x^{k+1},$ (21)
where $(s)_{k}=\frac{\Gamma(s+k)}{\Gamma(s)}=s(s+1)\cdots(s+k-1)$ if $k\geq 1$ and $(s)_{0}=1.$ We use now the well-known equivalence related to Euler summation (cf.[H1])
$\sum_{m\geq 0}b_{m}(\frac{x}{1+x})^{m+1}=\sum_{k\geq 0}a_{k}x^{k+1}\ \Leftrightarrow\ b_{m}=\sum_{k=0}^{m}\binom{m}{k}\ a_{k},$
which gives, for $P_{m}(s),$ the following explicit expression
$P_{m}(s)=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k+1}\ \frac{(s)_{k}}{k!}.$ (22)
We also have an integral expression for $P_{m}(s)$ when $0<\Re(s)<1$, since for $k\geq 0$, we can write
$\frac{(s)_{k}}{k!}=\frac{1}{\Gamma(s)\Gamma(1-s)}\int_{0}^{1}t^{s-1+k}(1-t)^{-s}dt,$
thus, by (22), we get for $0<\Re(s)<1$
$P_{m}(s)=2\frac{\sin(\pi s)}{\pi}\int_{0}^{1}t^{s-1}(1-t)^{-s}(1-2t)^{m}dt.$ (23)
This integral expression allows us to obtain the following lemma which will be useful later.
Lemma 1.
For $0<\Re(s)<1$, we have, for every $m\geq 1$
$|P_{m}(s)|\leq e^{\pi|Im(s)|}\Big(\frac{\Gamma(\Re(s))}{m^{\Re(s)}}+\frac{\Gamma(1-\Re(s))}{m^{1-\Re(s)}}\Big).$
Proof.
Using (23) we get
$|P_{m}(s)|$ $\leq$ $2\frac{|\sin(\pi s)|}{\pi}\int_{0}^{1}t^{\Re(s)-1}(1-t)^{-\Re(s)}|1-2t|^{m}dt,$
$\leq$ $2\frac{|\sin(\pi s)|}{\pi}\int_{-1}^{1}(1-u)^{\Re(s)-1}(1+u)^{-\Re(s)}|u|^{m}du,$
$\leq$ $e^{\pi|Im(s)|}\int_{-1}^{1}(1-u)^{\Re(s)-1}(1+u)^{-\Re(s)}|u|^{m}du.$
Let $a=\Re(s)$, this last integral is :
$I_{a}=\int_{0}^{1}(1-u)^{a-1}(1+u)^{-a}u^{m}du+\int_{-1}^{0}(1-u)^{a-1}(1+u)^{-a}(-u)^{m}du.$
Since $0<a<1$, we have
$I_{a}\leq\int_{0}^{1}(1-u)^{a-1}u^{m}du+\int_{-1}^{0}(1+u)^{-a}(-u)^{m}du$ $=$ $\int_{0}^{1}(1-u)^{a-1}u^{m}du+\int_{0}^{1}(1-u)^{-a}u^{m}du.$
Now if $0<\alpha<1$ we have, for every $x>0$
$\int_{0}^{1}{(1-u)^{\alpha-1}}u^{x}du=\int_{0}^{+\infty}{e^{-t}}{(1-e^{-t})^{\alpha-1}}e^{-xt}dt\leq\int_{0}^{+\infty}t^{\alpha-1}e^{-xt}dt=\Gamma(\alpha)x^{-\alpha},$
(the last inequality is justified by $1\leq\frac{t}{1-e^{-t}}\leq{e^{t}}$, for every $t>0$). Finally we obtain
$I_{a}\leq\frac{\Gamma(a)}{m^{a}}+\frac{\Gamma(1-a)}{m^{1-a}}.$
Remark 6.
It is easily verified (cf. [G-R]) that, for $m\geq 1$, we have $P_{m}(s)=2F(-m,s,1,2),$ where $F$ is the Gauss hypergeometric function
$F(a,b,c,z)=\sum_{k\geq 0}\frac{(a)_{k}(b)_{k}}{(c)_{k}}\frac{z^{k}}{k!},$
By classical properties of hypergeometric function (cf. [G-R]), we get the relations
$P_{m}(1-s)=(-1)^{m}P_{m}(s),$ (24)
$sP_{m}(s+1)-(s-1)P_{m}(s-1)=(2m+1)P_{m}(s).$ (25)
3.2 Orthogonality relations
Let $(\varphi_{m})_{m\geq 0}$ the sequence of functions
$\varphi_{m}:x\mapsto\sqrt{2}\ e^{-x}L_{m}(2x).$
It is well known (cf. [Leb]) that the sequence $(\varphi_{m})_{m\geq 0}$ is an orthonormal system in $(L^{2}(0,+\infty),dx)$. We verify as in the introduction that, by Mellin transform, we have for $\Re(s)>0$
$\mathcal{M}\varphi_{m}(s)=\int_{0}^{+\infty}\varphi_{m}(x)x^{s-1}\ dx=\frac{1}{\sqrt{2}}{\Gamma(s)}P_{m}(s).$ (26)
Thus the Mellin-Parseval formula (cf. [T]), gives
$\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\mathcal{M}(\varphi_{m})(\frac{1}{2}+it)}\ \mathcal{M}(\varphi_{n})(\frac{1}{2}+it)\ dt=\int_{0}^{+\infty}\overline{\varphi_{m}(x)}\varphi_{n}(x)dx=\delta_{(m,n)},$
with $\delta$ the usual Kronecker symbol. That is
$\frac{1}{4\pi}\int_{-\infty}^{+\infty}\overline{\Gamma(\frac{1}{2}+it)}\overline{P_{m}(\frac{1}{2}+it)}\ \Gamma(\frac{1}{2}+it)P_{n}(\frac{1}{2}+it)dt=\delta_{(m,n)}.$
Since
$\Gamma(\frac{1}{2}+it)\overline{\Gamma(\frac{1}{2}+it)}=\frac{\pi}{\cosh\pi t},$
then
$\frac{1}{4}\int_{-\infty}^{+\infty}\overline{P_{m}(\frac{1}{2}+it)}\ P_{n}(\frac{1}{2}+it)\frac{dt}{\cosh\pi t}=\delta_{(m,n)}.$
This proves that the sequence $(Q_{m})_{m\geq 0}$ of polynomials
$Q_{m}:t\mapsto\frac{1}{2}P_{m}(\frac{1}{2}+it)=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k}\ \frac{(\frac{1}{2}+it)_{k}}{k!}$ (27)
is an orthonormal system in $\big(L^{2}(-\infty,+\infty),\frac{dt}{\cosh(\pi t)}\big)$. It is also an orthonormal basis of this space since for $|a|<\pi$, we have
$\int_{-\infty}^{+\infty}e^{a|t|}\frac{dt}{\cosh(\pi t)}<+\infty$
By a classical result on orthogonal polynomials we deduce that, for every $m\geq 1$, the roots of $Q_{m}$ are real. This implies that the roots of $P_{m}$ are on the line $\{\Re(s)=\frac{1}{2}\}$.
For the first values of $m$ we get
$Q_{0}(t)$ $=$ $1,$
$Q_{1}(t)$ $=$ $i(-2t),$
$Q_{2}(t)$ $=$ $-2t^{2}+\frac{1}{2},$
$Q_{3}(t)$ $=$ $i(\frac{4}{3}t^{3}-\frac{5}{3}t),$
$Q_{4}(t)$ $=$ $\frac{2}{3}t^{4}-\frac{7}{3}t^{2}+\frac{3}{8}.$
Remark 7.
By (24) and (25) we have, for every $m\geq 0$, the relations
$\overline{Q_{m}(t)}=Q_{m}(-t)=(-1)^{m}Q_{m}(t),$
$\big(\frac{1}{2}+it\big)Q_{m}(t-i)+\big(\frac{1}{2}-it\big)Q_{m}(t+i)=(2m+1)Q_{m}(t).$
3.3 Expansion of $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}$
We now apply Ramanujan’s summation on both sides of (18). Using (20), we get
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}G_{s}(\frac{n-1}{n+1})=\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s).$
The following theorem says that, for $0<\Re(s)<1$, these last two sums can be inverted.
Theorem 6.
For $0<\Re(s)<1$ we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\sum_{m=0}^{+\infty}z_{m}P_{m}(s)\ \text{ where }\ z_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}.$
And we obtain the expansion
$\zeta(s)=\frac{1}{s-1}+\sum_{m=0}^{+\infty}z_{m}P_{m}(s).$ (28)
Proof.
By definition 2, we can write
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}=c_{m}+i\int_{0}^{+\infty}\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\frac{1}{e^{2\pi t}-1}\ dt$
with $c_{0}=1/4$ if $m=0$ and $c_{m}=0$ if $m\geq 1$. We will now prove that the series
$\sigma(s)=\sum_{m\geq 0}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s)$
is convergent for every $0<\Re(s)<1$, and that its sum is equal to
$\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s).$
Since $c_{0}P_{0}(s)=1/2$, then, by definition 2, the series $\sigma(s)$ is also
$\sigma(s)=\frac{1}{2}+i\sum_{m\geq 0}\int_{0}^{+\infty}\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\frac{1}{e^{2\pi t}-1}\ dt\ P_{m}(s).$ (29)
To prove the convergence of this series and invert signs $\sum$ and $\int$ we use the following lemma.
Lemma 2.
For $0<\Re(s)<1$, the series
$\sum_{m\geq 0}\Big|\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\ P_{m}(s)\Big|$ (30)
is convergent and
$\int_{0}^{+\infty}\sum_{m=0}^{+\infty}\Big|\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\ P_{m}(s)\Big|\ \frac{1}{e^{2\pi t}-1}dt<+\infty.$ (31)
Proof of the lemma.
We have
$\Big|\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big|\leq 2\Big|\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}\Big|=\frac{2}{\sqrt{t^{2}+4}}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}$
Using lemma 1, we see that the terms of the series (30) are, for $m\geq 1$, dominated by the terms
$\frac{4}{\sqrt{t^{2}+4}}e^{\pi|Im(s)|}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}\Big(\frac{\Gamma(\Re(s))}{m^{\Re(s)}}+\frac{\Gamma(1-\Re(s))}{m^{1-\Re(s)}}\Big)$
Consequently, the series (30) is convergent, and we have
$\sum_{m=0}^{+\infty}\Big|\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\ P_{m}(s)\Big|\ \leq M(t)$
with
$M(t)=\frac{4t}{4+t^{2}}+e^{\pi|Im(s)|}\frac{4}{\sqrt{t^{2}+4}}\sum_{m=1}^{+\infty}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}\Big(\frac{\Gamma(\Re(s))}{m^{\Re(s)}}+\frac{\Gamma(1-\Re(s))}{m^{1-\Re(s)}}\Big)$
It remains to prove that
$\int_{0}^{+\infty}M(t)\frac{1}{e^{2\pi t}-1}dt<+\infty.$
This follows from the convergence of the integral
$\int_{0}^{+\infty}\frac{4t}{4+t^{2}}\frac{1}{e^{2\pi t}-1}dt,$
and the fact that, for $0<\alpha<1$, we have
$\int_{0}^{+\infty}\frac{1}{\sqrt{t^{2}+4}}\sum_{m=1}^{+\infty}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}\frac{1}{m^{\alpha}}\frac{1}{e^{2\pi t}-1}dt<+\infty,$
as a consequence of the following inequalities
$\frac{1}{\sqrt{t^{2}+4}}\sum_{m=2}^{+\infty}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}\frac{1}{m^{\alpha}}\leq\frac{1}{\sqrt{t^{2}+4}}\sum_{m=1}^{+\infty}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}\leq\frac{t}{{t^{2}+4}}\Big(\frac{1}{1-\frac{t}{\sqrt{t^{2}+4}}}\Big)\leq\frac{t^{2}}{4}.$
End of the proof of the lemma.
By the preceding lemma the series $\sigma(s)$ is convergent for every $0<\Re(s)<1$, and we get
$\sum_{m=0}^{+\infty}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s)$
$\frac{1}{2}+i\int_{0}^{+\infty}\Big(\frac{1}{2+it}\sum_{m=0}^{+\infty}\big(\frac{it}{2+it}\big)^{m}P_{m}(s)-\frac{1}{2-it}\sum_{m=0}^{+\infty}\big(\frac{-it}{2-it}\big)^{m}P_{m}(s)\Big)\frac{1}{e^{2\pi t}-1}\ dt.$
Using (20) with $z=\frac{\pm it}{2\pm it}$, this last expression is equal to
$\frac{1}{2}+i\int_{0}^{+\infty}\Big(\big(\frac{1}{1+it}\big)^{s}-\big(\frac{1}{1-it}\big)^{s}\Big)\frac{1}{e^{2\pi t}-1}\ dt,$
which is by definition 2, also equal to
$\frac{1}{2}+\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}-\frac{1}{2}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}.$
The relation (28) follows immediately from (17).
3.4 Evaluation of the sum $z_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}$
If we apply the shift formula of corollary 1 to the function $f:x\mapsto\frac{1}{x}\big(1-\frac{2}{x}\big)^{m}$, we get
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}\Big(1-\frac{2}{n}\Big)^{m}-(-1)^{m}+y_{m}$
with
$y_{m}=\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}dx=\int_{1}^{2}\Big(1-\frac{2}{x}\Big)^{m}\frac{1}{x}dx.$
Now the binomial expansion
$\Big(1-\frac{2}{x}\Big)^{m}=\sum_{k=0}^{m}\binom{m}{k}(-2)^{k}\frac{1}{x^{k}}$
gives
$z_{m}=\sum_{k=0}^{m}\binom{m}{k}(-2)^{k}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{k+1}}-(-1)^{m}+\sum_{k=0}^{m}\binom{m}{k}(-2)^{k}\int_{1}^{2}\frac{1}{x^{k+1}}dx.$
Thus we obtain $z_{m}=x_{m}+y_{m}$, where $x_{0}=\gamma-1,\ \ y_{0}=\log(2),$ and for $m\geq 1$
$x_{m}$ $=$ $\gamma+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\Big(\zeta(k+1)-\frac{1}{k}\Big)-(-1)^{m},$ (32)
$y_{m}$ $=$ $\log(2)+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\frac{1}{k}\Big(1-\frac{1}{2^{k}}\Big).$
For the first values of $m$ we get
$x_{0}$ $=$ $\gamma-1,$
$x_{1}$ $=$ $\gamma-2\zeta(2)+3,$
$x_{2}$ $=$ $\gamma-4\zeta(2)+4\zeta(3)+1,$
$x_{3}$ $=$ $\gamma-6\zeta(2)+12\zeta(3)-8\zeta(4)+11/3,$
$x_{4}$ $=$ $\gamma-8\zeta(2)+24\zeta(3)-32\zeta(4)+16\zeta(5)+5/3.$
Theorem 7.
For $0<\Re(s)<1$, we have
$\zeta(s)=\frac{1}{s-1}+\sum_{m=0}^{+\infty}z_{m}P_{m}(s),$ (33)
where
$z_{m}=\gamma+\log(2)+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\zeta(k+1)+(-1)^{m+1}+\sum_{k=1}^{m}\frac{1}{k}.$
Proof.
This follows from the expansion (28) of theorem 6, and the combinatorial identity
$\sum_{k=1}^{m}\binom{m}{k}\frac{(-1)^{k-1}}{k}=\sum_{k=1}^{m}\frac{1}{k}.$
3.5 An expansion of Zeta
The previous theorem will enable us to get an expansion of Zeta in the strip $\{0<\Re(s)<1\}$. Indeed, the polar part 1/(s1)1/(s-1) of (33) has the simple expansion given by the following lemma.
Lemma 3.
For $0<\Re(s)<1$ we have
$\frac{1}{s-1}=-\sum_{m=0}^{+\infty}y_{m}P_{m}(s).$ (34)
Proof.
For $0<\Re(s)<1$, we write
$\sum_{m=0}^{+\infty}y_{m}P_{m}(s)=\sum_{m=0}^{+\infty}\Big(\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}dx\Big)P_{m}(s)=\int_{0}^{1}\sum_{m=0}^{+\infty}\Big(\frac{x-1}{x+1}\Big)^{m}P_{m}(s)\frac{1}{x+1}dx.$ (35)
This interchanging of signs $\sum$ and $\int$ is justified by
$\sum_{m=0}^{+\infty}\int_{0}^{1}|\frac{x-1}{x+1}|^{m}\ |P_{m}(s)|\frac{1}{x+1}dx\leq\sum_{m=0}^{+\infty}\int_{0}^{1}\frac{1}{(x+1)^{m+1}}dx\ |P_{m}(s)|\leq\sum_{m=0}^{+\infty}\frac{1}{m}|P_{m}(s)|,$
and we immediately check, with the lemma 1, that the latter series is convergent.
The relations (18) and (20) allows us to write for every $x>0$
$\sum_{m=0}^{+\infty}\frac{1}{x+1}\Big(\frac{x-1}{x+1}\Big)^{m}P_{m}(s)=\frac{1}{x^{s}},$
and by (35) we have
$\sum_{m=0}^{+\infty}y_{m}P_{m}(s)=\int_{0}^{1}\frac{1}{x^{s}}dx=-\frac{1}{s-1}.$
Theorem 8.
For $0<\Re(s)<1$, we have
$\zeta(s)=\sum_{m=0}^{+\infty}x_{m}P_{m}(s)$ (36)
with
$x_{m}=\gamma+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\zeta(k+1)+r_{m}$ (37)
and $r_{m}$ are the rational numbers
$r_{m}=(-1)^{m+1}+\sum_{k=1}^{m}\frac{1+(-1)^{k-1}}{k}$
Proof.
This is an immediate consequence of the preceding lemma and theorem 7 and the combinatorial identity
$\sum_{k=1}^{m}\binom{m}{k}\frac{(-2)^{k-1}}{k}=\sum_{k=1}^{m}\frac{1+(-1)^{k-1}}{k}.$
3.6 An orthonormal expansion
Let $X(0)=\gamma-1$ and for $k\geq 1$
$X(k)=\zeta(k+1)-1-\frac{1}{k}.$
Then formula (32) can be written as a binomial transform
$x_{m}=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k}X(k).$
Let
$f_{0}(t)=\frac{1}{e^{t}-1}-\frac{1}{t},$
then we have for $k\geq 0$
$X(k)=\int_{0}^{+\infty}e^{-t}\frac{t^{k}}{k!}\ f_{0}(t)dt$
By definition of Laguerre polynomials, we obtain for the coefficients $x_{m}$ the integral expression
$x_{m}=\int_{0}^{+\infty}e^{-t}L_{m}(2t)f_{0}(t)dt=\frac{1}{\sqrt{2}}\int_{0}^{+\infty}\varphi_{m}(t)f_{0}(t)dt=\frac{1}{\sqrt{2}}(\varphi_{m},f_{0}).$ (38)
Theorem 9.
We have for any integer $m\geq 0$
$x_{m}=\frac{1}{2}\int_{-\infty}^{+\infty}\overline{Q_{m}(t)}\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}.$ (39)
Proof.
By Mellin-Parseval formula (cf. [T]), we have
$\frac{1}{\sqrt{2}}\int_{0}^{+\infty}{\varphi_{m}(t)}f_{0}(t)dt=\frac{1}{2\pi\sqrt{2}}\int_{-\infty}^{+\infty}\overline{\mathcal{M}(\varphi_{m})(\frac{1}{2}+it)}\ \mathcal{M}(f_{0})(\frac{1}{2}+it)\ dt$
Since (cf. [T]) we have $\mathcal{M}f(s)=\Gamma(s)\zeta(s)$ for $0<\Re(s)<1$, then, using (26), we get
$x_{m}=\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\Gamma(\frac{1}{2}+it)\frac{1}{2}P_{m}(\frac{1}{2}+it)}\Gamma((\frac{1}{2}+it))\frac{1}{2}\zeta(\frac{1}{2}+it)\ dt,$
since $Q_{m}(t)=\frac{1}{2}P_{m}(\frac{1}{2}+it)$, this is the integral expression (39).
The previous theorem means that the pointwise convergent expansion
$\zeta(\frac{1}{2}+it)=2\sum_{m=0}^{+\infty}x_{m}Q_{m}(t)$
is also an orthonormal expansion in the space $\big(L^{2}(-\infty,+\infty),\frac{dt}{\cosh(\pi t)}\big)$.
Then we have
$\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=4\sum_{m=0}^{+\infty}x_{m}^{2}.$ (40)
We give next a formal evaluation of this last sum using Ramanujan’s summation.
3.7 A generating function
Let $\mathcal{L}$ the Laplace transform
$\mathcal{L}f(z)=\int_{0}^{+\infty}e^{-xz}f(x)\ dx=(e^{-xz},f).$
By Paley-Wiener theorem (cf. [Ho], [P]) , the transformation $\frac{1}{\sqrt{2\pi}}\mathcal{L}$ maps isometrically the space $L^{2}(]0,+\infty[,dx)$ to the Hardy space $H^{2}(\mathbb{C}_{+})$ of analytic functions $g$ in the half-plane $\mathbb{C}_{+}=\{z|\Re(z>0\}$, such that : there exists $M_{g}>0$ with
$\int_{\mathbb{R}}|g(x+iy)|^{2}\ dy\leq M_{g}\text{ for all }x>0.$
Each such $g$ has non-tangential limits $g(iy)=\lim_{x\to 0}g(x+iy)$ at almost every point of the imaginary axis. The space $H^{2}(\mathbb{C}_{+})$ is an Hilbert space with the inner product
$(g,h)_{H^{2}}=\int_{\mathbb{R}}\overline{g(iy)}{h(iy)}dy,$
and the functions $\psi_{m}$ $(m\geq 0)$, form an orthonormal basis of $H^{2}(\mathbb{C}_{+})$. If $f\in L^{2}(]0,+\infty[,dx)$. From
$e^{-xz}=\sqrt{2}\ \sum_{k=0}^{+\infty}\varphi_{k}(x)\psi_{k}(z).$
we get
$\frac{1}{\sqrt{2\pi}}\mathcal{L}\varphi_{m}(z)=\frac{1}{\sqrt{2\pi}}(e^{-xz},\varphi_{m})=\frac{1}{\sqrt{\pi}}\psi_{m}(z)$
The system $(\frac{1}{\sqrt{\pi}}\psi_{m})_{m\geq 0}$ is an orthonormal basis of $H^{2}(\mathbb{C}_{+})$. From the expansion
$f_{0}=\sum_{m\geq 0}(\varphi_{m},f_{0})\varphi_{m}$
we get the expansion
$\frac{1}{\sqrt{2\pi}}\mathcal{L}f_{0}(z)=\sum_{m\geq 0}(\varphi_{m},f_{0})\frac{1}{\sqrt{\pi}}\psi_{m}(z).$
and, by (38) we deduce
$\mathcal{L}f_{0}(z)=2\sum_{m\geq 0}x_{m}\psi_{m}(z).$
Since it is well known (cf. [G-R]) that
$\int_{0}^{+\infty}e^{-xz}f_{0}(x)dx=\log(z)-\Psi(z)-\frac{1}{z}$
then we get
$2\sum_{m=0}^{+\infty}x_{m}\psi_{m}(z)=\log(z)-\Psi(z)-\frac{1}{z}.$ (41)
Equivalently this gives the generating function
$\sum_{m=0}^{+\infty}x_{m}u^{m}=\frac{1}{1-u}\Big(\log(\frac{1+u}{1-u})-\Psi(\frac{1+u}{1-u})\Big)-\frac{1}{1+u}.$ (42)
Remark 8.
Now, if we apply (41) with $z=1,2,3,\dots$ and take Ramanujan’s summation, we get
$2\sum_{m=0}^{+\infty}x_{m}\sum_{n\geq 1}^{\mathcal{R}}\psi_{m}(n)=2\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}x_{m}\psi_{m}(n)=\sum_{n\geq 1}^{\mathcal{R}}\log(n)-\sum_{n\geq 1}^{\mathcal{R}}\Psi(n)-\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}.$
Then, with the results of examples 4-6, this relation gives
$2\sum_{m=0}^{+\infty}x_{m}(x_{m}+y_{n})=\log(2\pi)-\gamma-\frac{3}{2}.$
Since $y_{m}=\int_{0}^{1}\psi_{m}(x)dx$ then by (41) we have
$2\sum_{m=0}^{+\infty}x_{m}y_{n}=\int_{0}^{1}2\sum_{m=0}^{+\infty}x_{m}\psi_{m}(x)dx=\int_{0}^{1}(\log(x)-\Psi(1+x))dx=-1.$
Finally we get
$\sum_{m=0}^{+\infty}x_{m}^{2}=\frac{1}{2}\log(2\pi)-\frac{1}{2}\gamma-\frac{1}{4}.$
Then (40) gives us the relation
$\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=2\log(2\pi)-2\gamma-1,$ (43)

### v2

@@
 Abstract
-After a brief introduction to Ramanujan’s method of summation, we give an expansion of the Riemann Zeta function in the critical strip as a convergent series $\sum_{m\geq 0}x_{m}P_{m}(s)$ where the functions $P_{m}$ are polynomials with their roots on the line $\{\Re(s)=1/2\}$, the coefficients $x_{m}$ being finite linear combinations of the Euler constant $\gamma$ and the values $\zeta(2),\zeta(3),\dots,\zeta(m+1).$
+We introduce a family of orthogonal polynomials $P_{m}(s)$, whose roots lie on the line $\{\Re(s)=1/2\}$, and we give an expansion of the Riemann zeta function in the critical strip as a convergent series $\sum_{m\geq 0}x_{m}P_{m}(s)$.
+The coefficients $x_{m}$of this expansion are expressed as linear combinations with rational coefficients of the Euler constant $\gamma$ and the values $\zeta(2),$ $\zeta(3),$ $\dots,$ $\zeta(m+1).$
 1 Introduction
-There are many polynomial expansions of the Riemann zeta function.
-Some of them (cf.[E-K-K], [F-V],[B],[M], [V]) are based on the Newton interpolation formula for functions related to Zeta.
-For example
+Some polynomial expansions of the Riemann zeta function (cf.[E-K-K], [F-V],[B],[M], [V]) are based on the Newton interpolation formula for functions related to Zeta. For example
 $\zeta(s)-\frac{1}{s-1}=\sum_{m=0}^{+\infty}(-1)^{m}c_{m}\binom{s-1}{m}$,
@@
 $c_{m}=\gamma+\sum_{k=1}^{m}\binom{m}{k}(-1)^{k}\big(\zeta(k+1)-\frac{1}{k}\big).$
-Others are linked to the well known expansion
+Others (cf.[E-G]) are linked to the well known expansion
 $\zeta(s)-\frac{1}{s-1}=\sum_{m=0}^{+\infty}(-1)^{m}\gamma_{m}\frac{(s-1)^{m}}{m!}$
-where $\gamma_{m}$ are the Stieltjes constants. This gives (cf.[E-G])
-$\zeta(s)-\frac{s}{s-1}=\sum_{m=0}^{+\infty}(-1)^{m}l_{m}(\frac{s-1}{s})^{m},$
-with
-$l_{m}=\sum_{k=1}^{m}\binom{m-1}{k-1}(-1)^{m-k}\frac{\gamma_{k}}{k!}.$
-and for $s=\frac{1}{2}+it$ $(t\in\mathbb{R}$) one obtain an expansion of $\zeta(\frac{1}{2}+it)$ on the orthonormal basis $(e_{m})$ of $L^{2}(\mathbb{R},\frac{dt}{2\pi(t^{2}+1/4)})$ given by
-$e_{m}(t)=\Big(\frac{1-2it}{1+2it}\Big)^{m}.$
-There is also (cf. [K]) an expansion of the function
-$\Xi(t)=\frac{s(s-1)}{2}\pi^{-s/2}\Gamma(s/2)\zeta(s),\ \ \ s=\frac{1}{2}+it,\ (t\in\mathbb{R}),$
-which takes the form
-$\Xi(t)=e^{\pi t/4}\sum_{m=0}^{+\infty}b_{m}p_{m}(t),$
-where $p_{m}(t)$ are polynomials related to the Mellin transform of odd Hermite functions.
-These polynomials are orthogonal in $(L^{2}(\mathbb{R},|\Gamma(3/4+it/2)|^{2}dt)$, but it seems that the coefficients $b_{m}$ does not have a simple form.
-The polynomial expansion we give in this paper is related to the Mellin transform of the functions $x\mapsto e^{-x}L_{m}(2x)$ where $(L_{m})_{m\geq 0}$ are the Laguerre polynomials defined by
+where $\gamma_{m}$ are the Stieltjes constants.
+The polynomial expansion of the Riemann zeta function that we give in this paper is related to the Mellin transform of the functions $x\mapsto e^{-x}L_{m}(2x)$ where $(L_{m})_{m\geq 0}$ is the sequence of Laguerre polynomials defined by
 $L_{m}(x)=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}\frac{x^{k}}{k!}.$
 The generating function of the Laguerre polynomials (cf. [Leb]) gives for $x\in[0,+\infty[$
-$e^{-xz}=\sum_{m=0}^{+\infty}2e^{-x}L_{m}(2x)\psi_{m}(z).$
-where $(\psi_{m})_{m\geq 0}$ the sequence of the functions defined in the half-plane $\{Re(z)>0\}$ by
-$\psi_{m}(z)=\frac{1}{1+z}\big(\frac{z-1}{z+1}\big)^{m}.$
+$e^{-xz}=\sum_{m=0}^{+\infty}2e^{-x}L_{m}(2x)\big(\frac{z-1}{z+1}\big)^{m}\frac{1}{1+z}.$
 By Mellin transform of the function $x\mapsto e^{-xz}$, we obtain for $\Re(z)>0$ and $\Re(s)>0$
-$\Gamma(s)z^{-s}=\sum_{m=0}^{+\infty}\Big(2\int_{0}^{+\infty}e^{-x}L_{m}(2x)x^{s-1}\ dx\Big)\psi_{m}(z),$
-Since we can write
-$2\int_{0}^{+\infty}e^{-x}L_{m}(2x)x^{s-1}\ dx={\Gamma(s)}P_{m}(s),$
-where $P_{m}$ is the polynomial
-$P_{m}(s)=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k+1}\ \frac{\Gamma(s+k)}{\Gamma(s)}\frac{1}{k!},$
-then we have, for $\Re(s)>0$, the expansion
-$z^{-s}=\sum_{m=0}^{+\infty}\psi_{m}(z)P_{m}(s).$
+$z^{-s}=\sum_{m=0}^{+\infty}\frac{1}{1+z}\big(\frac{z-1}{z+1}\big)^{m}P_{m}(s),$
+where $P_{m}$ is a polynomial defined by
+${\Gamma(s)}P_{m}(s)=2\int_{0}^{+\infty}e^{-x}L_{m}(2x)x^{s-1}\ dx.$
 Take $z=1,2,3,\dots$, then we get an infinite sequence of relations that it’s tempting to sum up, to obtain
-$\sum_{n\geq 1}\frac{1}{n^{s}}=\sum_{m=0}^{+\infty}\Big(\sum_{n\geq 1}\psi_{m}(n)\Big)P_{m}(s).$
-In section 3, we prove that, for $0<\Re(s)<1$, this last formula is correct if we replace the divergent series $\sum_{n\geq 1}$ by their Ramanujan’s summations $\sum_{n\geq 1}^{\mathcal{R}}$, and then we have
+$\sum_{n\geq 1}\frac{1}{n^{s}}=\sum_{m=0}^{+\infty}\Big(\sum_{n\geq 1}\frac{1}{1+n}\big(\frac{n-1}{n+1}\big)^{m}\Big)P_{m}(s).$
+We prove that, for $0<\Re(s)<1$, this last formula is correct if we replace the divergent series $\sum_{n\geq 1}$ by their Ramanujan’s summations $\sum_{n\geq 1}^{\mathcal{R}}$, and then we have
 $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\sum_{m=0}^{+\infty}z_{m}P_{m}(s)\ \text{ where }\ z_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}.$
-In section 2 we give a short introduction to Ramanujan’s summation of divergent series. Since we will see that, for $s\neq 1$, we have
+We give in section 2, a short introduction to Ramanujan’s summation of divergent series. Since we will see that, for $s\neq 1$, we have
 $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\zeta(s)-\frac{1}{s-1},$
@@
 $x_{m}=\gamma+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\zeta(k+1)+r_{m}$
-and $r_{m}$ are rational numbers explicitely given. For $s=\frac{1}{2}+it$ with $t\in\mathbb{R}$, the preceding relation gives the expansion
+and $r_{m}$ are the rational numbers
+$r_{m}=(-1)^{m+1}+\sum_{k=1}^{m}\frac{1+(-1)^{k-1}}{k}.$
+We prove that the roots of $P_{m}$ are on the line $\{\Re(s)=\frac{1}{2}\}$. This is a consequence of the fact that the polynomials
+$Q_{m}:t\mapsto\frac{1}{2}P_{m}(\frac{1}{2}+it)$
+form a complete orthonormal system in $L^{2}\big(\mathbb{R},\frac{dt}{\cosh(\pi t)}\big)$. The expansion
 $\zeta(\frac{1}{2}+it)=2\sum_{m=0}^{+\infty}x_{m}Q_{m}(t)$
-where the polynomials
-$Q_{m}:t\mapsto\frac{1}{2}P_{m}(\frac{1}{2}+it)$
-form complete orthonormal system in $\big(L^{2}(-\infty,+\infty),\frac{dt}{\cosh(\pi t)}\big)$.
-Thus, for every $m\geq 1$, the roots of $Q_{m}$ are real, and this implies that the roots of $P_{m}$ are on the line $\{\Re(s)=\frac{1}{2}\}$.
+is an orthonormal expansion in the space $L^{2}\big(\mathbb{R},\frac{dt}{\cosh(\pi t)}\big)$. Also by a formal calculation, we get
+$\frac{1}{2}\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=\log(2\pi)-\gamma-\frac{1}{2}.$
 2 Ramanujan’s summation
 2.1 Ramanujan’s constant of a series
-Let a function $f\in C^{1}(]0,+\infty[)$, the first order Euler-MacLaurin formula says that, for every integer $n\geq 1$
-$\sum_{k=1}^{n}f(k)=\int_{1}^{n}f(x)\,dx+\frac{1}{2}\big(f(n)+f(1)\big)+\int_{1}^{n}\big(x-[x]-\frac{1}{2}\big)f^{\prime}(x)\ dx$}, (1)
-where $[x]$ denote the integer part of $x$. For $x\geq 1$, let
-$b(x)=x-[x]-\frac{1}{2}$
-and let’s assume that the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\ dx$ is convergent. The previous formula gives
+Let a function $f\in C^{1}(]0,+\infty[)$, then for every integer $n\geq 1$, we have the first order Euler-MacLaurin formula
+$\sum_{k=1}^{n}f(k)=\int_{1}^{n}f(x)\,dx+\frac{1}{2}\big(f(n)+f(1)\big)+\int_{1}^{n}b(x)f^{\prime}(x)\ dx$}, (1)
+where
+b(x)=x[x]12 and [x] denote the integer part of x.b(x)=x-[x]-$\frac${1}{2}\ $\text${ and }\ [x]\ $\text${ denote the integer part of }\ x.
+Let’s assume that the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\ dx$ is convergent. The previous formula gives
 $\sum_{k=1}^{n}f(k)=\int_{1}^{n}f(x)\,dx+C(f)+\frac{1}{2}f(n)-\int_{n}^{+\infty}b(x)f^{\prime}(x)\ dx$}, (2)
@@
 $C(f)=\sum_{k=1}^{+\infty}f(k)-\int_{1}^{+\infty}f(x)\ dx.$
-We observe that $C(f)$ is also well defined for some divergent series. We have immediately $C(1)=1/2.$ And applying (2) to the function $f:x\mapsto 1/x$, we have for every integer $n\geq 1$
+We observe that $C(f)$ is also well defined for some divergent series. For example, we have immediately $C(1)=1/2,$ and applying (2) to the function $f:x\mapsto 1/x$, we have for any integer $n\geq 1$
 $\sum_{k=1}^{n}\frac{1}{k}=\log n+C(f)+\frac{1}{2n}+\int_{n}^{+\infty}b(x)\frac{1}{x^{2}}\ dx,$
-which gives that the Ramanujan’s constant of the series $\sum_{n\geq 1}\frac{1}{n}$ is the Euler constant :
-$C(f)=\gamma=\lim_{n\to+\infty}\Big(\sum_{k=1}^{n}\frac{1}{k}-\log n\Big).$
+and we see that the Ramanujan’s constant of the series $\sum_{n\geq 1}\frac{1}{n}$ is the Euler constant :
+$\gamma=\lim_{n\to+\infty}\Big(\sum_{k=1}^{n}\frac{1}{k}-\log n\Big).$
 2.2 Plana’s formula
@@
 Theorem 1.
-Let $f$ be an analytic function in the half-plane $\{\Re(z)>0\}$ such that there is a constant ${a}<2\pi$ and, for any interval $c,d]\subset]0,+\infty[$, a constant $C_{c,d}>0$ with
-$\left|f(x+it)\right|\leq C_{c,d}\ e^{a|t|},\ \text{ for }x\in[c,d]\ \text{ and all }t\in\mathbb{R}.$
+Let $f$ be an analytic function in the half-plane $\{\Re(z)>0\}$ such that there is a constant ${a}<2\pi$ and, for any interval $c,d]\subset]0,+\infty[$, a constant $M>0$ with
+$\left|f(x+it)\right|\leq M\ e^{a|t|},\ \text{ for }x\in[c,d]\ \text{ and }t\in\mathbb{R}.$
 Then we have Plana’s formula : for any integer $n\geq 1$,
@@
 $+i\!\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\,\ dt-i\!\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\,\ dt\,.$ (3)
-If we now compare Euler-MacLaurin formula (1) and Plana’s (3), then we get, for $n\geq 1$,
+If we now compare Euler-MacLaurin formula (1) and Plana’s formula (3), then we get, for $n\geq 1$,
 $\int_{1}^{n}b(x){f^{\prime}(x)}\ dx=i\!\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\ dt-i\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\,\ dt\,.$
-If the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\,dx$ is convergent and if
+Then if the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\,dx$ is convergent and if
 $\lim_{n\to+\infty}\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\ dt=0,$
-then we have
+we have
 $\int_{1}^{+\infty}b(t){f^{\prime}(t)}\ dt=i\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\,\ dt.$
-Thus, in this case the Ramanujan’s constant $C(f)$, previously defined by (2), is also
+Thus, in this case the Ramanujan’s constant $C(f)$, previously defined in (2), is also
 $C(f)=\frac{1}{2}f(1)+i\int_{0}^{+\infty}\frac{{f(1+it)-f(1-it)}}{e^{2\pi t}-1}\,\ dt.$ (4)
-Take now this last formula as the definition of Ramanujan’s constant $C(f)$ for a function $f$ verifying the hypotheses of theorem 1.
 Remark 1.
@@
 Definition 1.
-Let ${\alpha}>0$ and ${\mathcal{E}}^{\alpha}$ the space of analytic functions $f$ in the half plane $\{\Re(z)>0\}$ such that there is some constants $\emph{a}<\alpha$ and $\emph{b}>0$ with :
+Let $\mathcal{E}$} the space of analytic functions $f$ in the half plane $\{\Re(z)>0\}$ such that there is some constants $\emph{a}<\pi$ and $\emph{b}>0$ with :
 for every ${\varepsilon}>0$, there is a constant $C_{\varepsilon}$ such that
@@
 $|f_{s}(z)|\;\;=\,\;e^{\sigma\Re(z)-\tau\Im(z)}.$
-Thus $f_{s}\in\mathcal{E}^{\alpha}$ if and only if $\alpha<\Im(s)<\alpha$.
-Properties
-a) If $f\in\mathcal{E}^{\alpha}$ and $g\in\mathcal{E}^{\beta}$, alors $fg\in E^{\alpha+\beta}$.
-b) If $f\in\mathcal{E}^{\alpha}$ and if $F(z)=\int_{1}^{z}f(u)du$ then $F\in\mathcal{E}^{\alpha}$.
+Thus $f_{s}\in\mathcal{E}$ if and only if $\pi<\Im(s)<\pi$.
 An important theorem for the rest of this section is the following.
@@
 Carlson’s theorem (cf. [H2])
-If $f\in\mathcal{E}^{\pi}$ is such that $f(n)=0$ for all $n\;=\;1,\,2,\,\ldots$, then $f\equiv 0$.
-Thus a function $f\in\mathcal{E}^{\pi}$ is only determined by the values $f(1),f(2),f(3),\dots$.
+If $f\in\mathcal{E}$ is such that $f(n)=0$ for all $n\;=\;1,\,2,\,\ldots$, then $f\equiv 0$.
+Thus a function $f\in\mathcal{E}$ is only determined by the values $f(1),f(2),f(3),\dots$.
 Remark 2.
-The function $g\;:\;z\mapsto\sin(\pi z)$ is such that $g(n)=0$ for every integer $n\geq 1$}, but $g\notin\mathcal{E}^{\pi}$ since for $x\in\mathbb{R}$ and $t\in\mathbb{R}$
+The function $g\;:\;z\mapsto\sin(\pi z)$ is such that $g(n)=0$ for every integer $n\geq 1$}, but $g\notin\mathcal{E}$ since for $x\in\mathbb{R}$ and $t\in\mathbb{R}$
 $|g(x+it)|\sim\frac{1}{2}e^{\pi|t|}\text{ when }t\to\infty$
@@
 Let $f$ and $g$ two functions verifying the hypothesis of theorem 1.
-By Carlson’s theorem we know that if $f\in\mathcal{E}^{\pi}$ and $g\in\mathcal{E}^{\pi}$ verify f(n)=g(n){f(n)=g(n)} for $n=1,2,3,\dots$, then for x>0{x>0} and $t\in\mathbb{R}$, we have
+By Carlson’s theorem we know that if $f\in\mathcal{E}$ and $g\in\mathcal{E}$ verify f(n)=g(n){f(n)=g(n)} for $n=1,2,3,\dots$, then for x>0{x>0} and $t\in\mathbb{R}$, we have
 f(x+it)f(xit)=g(x+it)g(xit),{f(x+it)-f(x-it)}={g(x+it)-g(x-it)},
@@
 Definition 2.
-For $f\in\mathcal{E}^{\pi}$, we define the Ramanujan’s summation of the series $\sum_{n\geq 1}f(n)$ by
-$\sum_{n\geq 1}^{\mathcal{R}}f(n)=\frac{f(1)}{2}+i\!\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\,\ dt\,.$ (6)
+For $f\in\mathcal{E}$, we define the Ramanujan’s summation of the series $\sum_{n\geq 1}f(n)$ by
+$\sum_{n\geq 1}^{\mathcal{R}}f(n)=C(f)=\frac{f(1)}{2}+i\!\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\,\ dt\,.$ (6)
 It is clear that $\sum_{n\geq 1}^{\mathcal{R}}f(n)$ depends linearly on $f$.
 Remark 3.
-Let $f\in\mathcal{E}^{\pi}$ then by (3), we have for every integer $n\geq 1$
+Let $f\in\mathcal{E}$ then by Plana’s formula (3), we have for every integer $n\geq 1$
 $\sum_{k=1}^{n}f(k)=\!\int_{1}^{n}f(u)\ du+\frac{f(n)}{2}+\sum_{n\geq 1}^{\mathcal{R}}f(n)-i\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\,\ dt\,.$ (7)
@@
 $\lim_{n\to+\infty}\int_{0}^{+\infty}\frac{f(n+it)-f(n-it)}{e^{2\pi t}-1}\,\ dt\,=0,$
-and if in addition, the integral $\int_{1}^{+\infty}f(u)\ du$ is convergent then the series $\sum_{n\geq 1}f(n)$ is convergent. In this case we have by (7)
+and if in addition, the integral $\int_{1}^{+\infty}f(u)\ du$ is convergent then the series $\sum_{n\geq 1}f(n)$ is convergent. In this case (7) gives the relation
 $\sum_{n\geq 1}^{\mathcal{R}}f(n)\;\;=\sum_{n=1}^{+\infty}f(n)-\!\!\int_{1}^{+\infty}\!\!f(x)\ dx.$
 Example 2.
-For $f:x\mapsto\frac{1}{x^{s}}$, with $\Re(s)>1$, we get
+Let $f:x\mapsto\frac{1}{x^{s}}$, then for $\Re(s)>1$, we get
 $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}\;\;=\sum_{n=1}^{+\infty}\frac{1}{n^{s}}-\!\!\int_{1}^{+\infty}\!\!\frac{1}{x^{s}}\ dx,$
@@
 2.5 Interpolation of partial sums
-To evaluate Ramanujan’s summation of a divergent series $\sum_{n\geq 1}f(n)$, a simpler way than the formula (6) can be obtained using an interpolation function of partial sums of this series .
+To evaluate Ramanujan’s summation of a divergent series $\sum_{n\geq 1}f(n)$, a simpler way than the formula (6) can be obtained using an interpolation function of the partial sums of this series.
 In chapter VI of his Notebooks, Ramanujan uses, without any definition, such a function $\varphi$ which satisfies
@@
 Proof.
-a) Existence. Since $f\in\mathcal{E}^{\pi}$ then the function $\varphi_{f}$, defined for $\Re(z)>0$, by
+a) Existence. Since $f\in\mathcal{E}$ then the function $\varphi_{f}$, defined for $\Re(z)>0$, by
 $\varphi_{f}(z)=\int_{1}^{z}f(u)\ du+\frac{f(z)}{2}+\sum_{n\geq 1}^{\mathcal{R}}f(n)-i\int_{0}^{+\infty}\frac{f(z+it)-f(z-it)}{e^{2\pi t}-1}\ dt,$ (10)
-is also in $\mathcal{E}^{\pi}$. By (3) we have $\varphi_{f}(n)=\sum_{k=1}^{n}f(k)$ for every integer $n\geq 1$, thus the function $\varphi_{f}$ verifies
+is also in $\mathcal{E}$. By (3) we have $\varphi_{f}(n)=\sum_{k=1}^{n}f(k)$ for every integer $n\geq 1$, thus the function $\varphi_{f}$ verifies
 $\varphi_{f}(n+1)-\varphi_{f}(n)\,=f(n+1),\text{ for }n=1,2,3,\dots.$
-Since the functions $z\mapsto\varphi_{f}(z+1)-\varphi_{f}(z)$ and $z\mapsto f(z+1)$ are in $\mathcal{E}^{\pi}$, then by Carlson’s theorem, we get
+Since the functions $z\mapsto\varphi_{f}(z+1)-\varphi_{f}(z)$ and $z\mapsto f(z+1)$ are in $\mathcal{E}$, then by Carlson’s theorem, we get
 $\varphi_{f}(z+1)-\varphi_{f}(z)=f(z+1),\quad\!\!\text{ for all }z\text{ such that }\Re(z)>0.\$(11)
-The condition $\varphi_{f}(1)=f(1)$ is simply a consequence of (6) and (10).
-b) Uniqueness. If $\varphi\in\mathcal{E}^{\pi}$ is such that
+The condition $\varphi_{f}(1)=f(1)$ is simply a consequence of (10) and definition (6).
+b) Uniqueness. If $\varphi\in\mathcal{E}$ is such that
 $\varphi(x+1)-\varphi(x)=f(x+1)\ \text{ for }x>0,\ \ \text{ with }\varphi(1)=f(1),$
@@
 thus $\varphi(n)=\sum_{k=1}^{n}f(k)=\varphi_{f}(n)$ for every integer $n\geq 1$.
-Since $\varphi$ and $\varphi_{f}$ are in the space $\mathcal{E}^{\pi}$, then by Carlson’s theorem we get $\varphi=\varphi_{f}.$
+Since $\varphi$ and $\varphi_{f}$ are in the space $\mathcal{E}$, then by Carlson’s theorem we get $\varphi=\varphi_{f}.$
 Remark 4.
-The preceding proof asserts that if $f\in\mathcal{E}^{\pi}$ then
+The preceding proof asserts that if $f\in\mathcal{E}$ then the function $\varphi_{f}$ given by (10) satisfies
 $\varphi_{f}(n)=\sum_{k=1}^{n}f(k)\text{ for every integer }n\geq 1.$
-The function $\varphi_{f}$ given by (10) is the unique function in $\mathcal{E}^{\pi}$ wich interpolates the partial sums of the series $\sum_{n\geq 1}f(n)$. Note that, by (11), we have for $\Re(z)>0$
+Thus this function $\varphi_{f}$ is the unique function in $\mathcal{E}$ which interpolates the partial sums of the series $\sum_{n\geq 1}f(n)$. Note that, by (11), we have for $\Re(z)>0$
 $\varphi_{f}(z)=\varphi_{f}(z+1)-f(z+1),$
@@
 Example 3.
-Let $f:z\mapsto 1/z$, then $\varphi_{f}$ is related to digamma function which is defined, for $\Re(z)>0$, by
+Let the function $f:z\mapsto 1/z$. We will ses that $\varphi_{f}$ is related to digamma function which is defined, for $\Re(z)>0$, by
 $\Psi(z)=\frac{\Gamma^{\prime}(z)}{\Gamma(z)}.$
-This function is analytic in $\{\Re(z)>0\}$, and by logarithmic differentiation of the classical product formula for $\Gamma$ (cf. [G-R]) we get for $\Re(z)>0$
-$\Psi(z)=-\frac{1}{z}-\gamma+\sum_{n=1}^{+\infty}\big(\frac{1}{n}-\frac{1}{n+z}\big)=-\frac{1}{z}-\gamma+\sum_{n=1}^{+\infty}\frac{z}{n(n+z)}$ (12)
-Then it is easily verified that $\Psi\in\mathcal{E}^{\pi}$ and that
-$\Psi(z+2)\;-\;\Psi(z+1)=\frac{1}{z+1}$
+The function $\Psi$ is analytic in $\{\Re(z)>0\}$, and by logarithmic differentiation of the classical product formula for $\Gamma$ (cf. [G-R]), we get for $\Re(z)>0$
+$\Psi(z)=-\frac{1}{z}-\gamma+\sum_{n=1}^{+\infty}\big(\frac{1}{n}-\frac{1}{n+z}\big)=-\frac{1}{z}-\gamma+\sum_{n=1}^{+\infty}\frac{z}{n(n+z)}.$
+Thus it is easily verified that $\Psi\in\mathcal{E}$ and that
+$\Psi(z+2)\;-\;\Psi(z+1)=\frac{1}{z+1}.$
 Since $\Psi(1)=-\gamma$, then, by the previous theorem, we get for $\Re(z)>0$
 $\varphi_{f}(z)=\Psi(z+1)+\gamma.$
-A very useful formula for Ramanujan’s summation is provided by the following theorem.
 Theorem 4.
-Let $f\in\mathcal{E}^{\pi}$, and let $\varphi\in\mathcal{E}^{\pi}$ be a solution of
-$\varphi(x+1)-\varphi(x)$ $=$ $f(x+1),\ \text{ for all }\ x>0,$
-$\varphi(1)$ $=$ $f(1).$
+Let $f\in\mathcal{E}$, and let $\varphi\in\mathcal{E}$ be a solution of
+$\varphi(x+1)-\varphi(x)=f(x+1),\ \text{ for all }\ x>0,\ \text{ and }\ \varphi(1)=f(1).$
 Then we have
@@
 By theorem 3, we have $\varphi=\varphi_{f}$. Let $F(z)=\int_{1}^{z}f(u)\ du$, then integration of (10) gives
-$\int_{1}^{2}\varphi_{f}(x)\ dx\;\;=\int_{1}^{2}F(u)\ du\,+\,\frac{1}{2}\int_{1}^{2}f(x)\ dx\,+\sum_{n\geq 1}^{\mathcal{R}}f(n)-iD(f),$ (13)
+$\int_{1}^{2}\varphi_{f}(x)\ dx\;\;=\int_{1}^{2}F(u)\ du\,+\,\frac{1}{2}\int_{1}^{2}f(x)\ dx\,+\sum_{n\geq 1}^{\mathcal{R}}f(n)-iD(f),$
 where
@@
 $D(f)=\int_{0}^{+\infty}\frac{F(2+it)-F(2-it)}{e^{2\pi t}-1}\ dt-\!\int_{0}^{+\infty}\frac{F(1+it)-F(1-it)}{e^{2\pi t}-1}\ dt.$
-If we apply Plana’s formula to the function $F$, with $n=2$, then we get
-$-\big(\frac{1}{2}F(1)+\frac{1}{2}F(2)\big)+\!\int_{1}^{2}F(t)dt$ $=i\int_{0}^{+\infty}\frac{F(2+it)-F(2-it)}{e^{2\pi t}-1}dt$
-$-i\int_{0}^{+\infty}\frac{F(1+it)-F(1-it)}{e^{2\pi t}-1}dt,$
-which is
+If we apply Plana’s formula (3) to the function $F$, in the case $n=2$, then we get
+$-\big(\frac{1}{2}F(1)+\frac{1}{2}F(2)\big)+\!\int_{1}^{2}F(t)\ dt=i\int_{0}^{+\infty}\frac{F(2+it)-F(2-it)}{e^{2\pi t}-1}\ dt-i\int_{0}^{+\infty}\frac{F(1+it)-F(1-it)}{e^{2\pi t}-1}\ dt,$
+that is
 $iD(f)=-\frac{F(2)}{2}+\int_{1}^{2}F(t)\ dt.$
-Then (13) becomes
+Then we get
 $\int_{1}^{2}\varphi_{f}(x)\ dx$ $=$ $\frac{1}{2}\int_{1}^{2}f(x)\ dx+\sum_{n\geq 1}^{\mathcal{R}}f(n)+\frac{F(2)}{2}$
-$=$ $\int_{1}^{2}f(x)\ dx+\sum_{n\geq 1}^{\mathcal{R}}f(n).$
+$=$ $\int_{1}^{2}f(x)\ dx+\sum_{n\geq 1}^{\mathcal{R}}f(n),$
 from which we deduce that
@@
 Corollary 1.
-Let $f\in\mathcal{E}^{\pi}$, then the function $x\mapsto f(x+1)$ belongs to $\mathcal{E}^{\pi}$ and we have the shift formula
-$\sum_{n\geq 1}^{\mathcal{R}}f(n+1)\;\;=\sum_{n\geq 1}^{\mathcal{R}}f(n)-f(1)+\int_{0}^{1}f(x+1)\ dx.$
+Let $f\in\mathcal{E}$, then the function $x\mapsto f(x+1)$ belongs to $\mathcal{E}$ and we have the shift formula
+$\sum_{n\geq 1}^{\mathcal{R}}f(n+1)\;\;=\sum_{n\geq 1}^{\mathcal{R}}f(n)-f(1)+\int_{0}^{1}f(x+1)\ dx.$ (12)
 Proof.
@@
 $\varphi_{g}(x)=\varphi_{f}(x+1)-f(1).$
-Thus we have
+Thus we obtain
 $\int_{0}^{1}\varphi_{g}(x)dx=\int_{0}^{1}\varphi_{f}(x+1)\ dx-f(1)=\int_{0}^{1}\varphi_{f}(x)\ \ dx+\int_{0}^{1}f(x+1)\ dx-f(1),$
@@
 Example 4.
-Let $f:x\mapsto 1/x$. By example 3, we have $\varphi_{f}(x)=\Psi(x+1)+\gamma,$ thus we get
+Let $f:z\mapsto 1/z$. By example 3, we have $\varphi_{f}(x)=\Psi(x+1)+\gamma,$ thus we get
 $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}=\int_{0}^{1}(\Psi(x+1)+\gamma)\ \ dx=\int_{0}^{1}\frac{\Gamma^{\prime}}{\Gamma}(x+1)\ dx+\gamma,$
@@
 $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}=\gamma,$ (13)
-Observe that, by corollary 1, we have
+Observe that, by the shift formula (12), we have
 $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}=\gamma-1+\log(2).$
 Example 5.
-Since, fro any $x>0$, we have
-$\log(\Gamma(x+2))-\log(\Gamma(x+1))=\log(x+1)$
+Let $\log$ the principal determination of logarithm. Since, for any $x>0$, we have
+$\log(\Gamma(x+2))-\log(\Gamma(x+1))=\log(x+1),$
 and $\log(\Gamma(2))=\log(1)$, we deduce that $\varphi_{\log}(x)=\log(\Gamma(x+1))$. Now if we use the Raabe’s formula, that is :
@@
 then we get
-$\sum_{n\geq 1}^{\mathcal{R}}\log(n)=\log(\sqrt{2\pi})-1.$
+$\sum_{n\geq 1}^{\mathcal{R}}\log(n)=\log(\sqrt{2\pi})-1.$ (14)
 Example 6.
-Let $f(x)=\Psi(x)$ and $\varphi(x)=x\Psi(x)-x+1$. We have $f(1)=\Psi(1)=\varphi(1)$ and
+Let $f=\Psi$ and let $\varphi:x\mapsto x\Psi(x)-x+1$. We have $f(1)=\Psi(1)=\varphi(1)$ and
 $\varphi(x+1)-\varphi(x)=\Psi(x+1)=f(x+1).$
-Thus, by the previous theorem, we get
+Thus, by the previous theorem, we have
 $\sum_{n\geq 1}^{\mathcal{R}}\Psi(n)=\int_{0}^{1}x\Psi(x)\ dx+\frac{1}{2}=-\int_{0}^{1}\log(\Gamma(x))\ dx+\frac{1}{2},$
-then by Raabe’s formula
-$\sum_{n\geq 1}^{\mathcal{R}}\Psi(n)=-\log(\sqrt{2\pi})+\frac{1}{2}.$
+then by Raabe’s formula, we get
+$\sum_{n\geq 1}^{\mathcal{R}}\Psi(n)=-\log(\sqrt{2\pi})+\frac{1}{2}.$ (15)
 Example 7.
-We can evaluate $\sum_{n\geq 1}^{\mathcal{R}}n^{k}$ for $k=1,2,3,\dots$ by means of Bernoulli numbers. Let the Bernoulli polynomials $B_{n}(x)$ be defined by $B_{0}(x)=1$ and, for $k\geq 1,$
-$B^{\prime}_{k}(x)=nB_{k-1}(x)\ \text{ with }\ \int_{0}^{1}B_{k}(x)dx=0.$
-The Bernoulli numbers are the numbers $B_{k}=B_{k}(0)$. These polynomials verify (cf. [G-R])
+We now evaluate $\sum_{n\geq 1}^{\mathcal{R}}n^{k}$ for any $k$ integer $\geq 1$. Let the Bernoulli polynomials $B_{k}(x)$ be defined by $B_{0}(x)=1$ and, for $k\geq 1,$
+$B^{\prime}_{k}(x)=kB_{k-1}(x)\ \text{ with }\ \int_{0}^{1}B_{k}(x)\ dx=0.$
+The Bernoulli numbers are the numbers $B_{k}=B_{k}(0)$. For $k\geq 1$, we have $B_{k+1}(1)=B_{k+1}$. These polynomials verify (cf. [G-R])
 $B_{k+1}\big((x+1)+1\big)-B_{k+1}(x+1)=(k+1)(x+1)^{k}.$
-For $k\geq 1$, we have $B_{k+1}(1)=B_{k+1}$. Let the function $f:x\mapsto x^{k}$ with an integer $k\geq 1$, then by theorem 3, we obtain
+Let the function $f:x\mapsto x^{k}$ with $k$ integer $\geq 1$. Then by theorem 3, we obtain
 $\varphi_{f}(x)=\frac{B_{k+1}(x+1)}{k+1}-\frac{B_{k+1}}{k+1}.$
@@
 Note also that
-$\sum_{n\geq 1}^{\mathcal{R}}n^{0}=\frac{1}{2}=\zeta(0)+1,$
-and by (9) we have in the half-plane $\{\Re(s)>1\}$
-$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\zeta(s)-\frac{1}{s-1}.$ (16)
-All these formulae are special cases the following general result.
+$\sum_{n\geq 1}^{\mathcal{R}}n^{0}=\frac{1}{2}=\zeta(0)+1.$
+These formulae are extensions of formula (8) and are special cases the following general result.
 Theorem 5.
@@
 is analytic on $\mathbb{C}$ and we have for every $s\in\mathbb{C}\setminus\{1\}$
-$\zeta(s)=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}+\frac{1}{s-1}$ (17)
+$\zeta(s)=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}+\frac{1}{s-1}.$ (17)
 Proof.
-First, it is easily seen that, for every $s\in\mathbb{C}$, the function $z\mapsto 1/{z^{s}}$ is in the space $\mathcal{E}^{\alpha}$, for all $\alpha>0.$ Since, if $s=\sigma+i\tau$ ($\sigma\in\mathbb{R},\ \tau\in\mathbb{R}$), we can write
+Let $s=\sigma+i\tau$ with $\sigma\in\mathbb{R},\ \tau\in\mathbb{R}$, and $z=x+it$ with $x>\varepsilon>0$, $t\in\mathbb{R}$. We have
 $\Big|\frac{1}{z^{s}}\Big|=\Big|\frac{1}{(x+it)^{s}}\Big|=\Big|e^{-s\log(x+it)}\Big|=e^{-{\sigma}\log\sqrt{x^{2}+t^{2}}+\tau\arctan(\frac{t}{x})}\leq e^{|\tau|\pi/2}(x^{2}+t^{2})^{-\sigma/2}$
-where $z=x+it$, with $x>\varepsilon>0$ and $t\in\mathbb{R}$. Thus the sum $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}$ is well defined and, by (6), we have
-$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\;\;\frac{1}{2}+i\int_{0}^{+\infty}\Big({\frac{1}{(1+it)^{s}}-\frac{1}{(1-it)^{s}}}\big)\frac{1}{e^{2\pi t}-1}\,\ dt\,$
+Thus for every $s\in\mathbb{C}$, the function $z\mapsto 1/{z^{s}}$ is in the space $\mathcal{E}$ and the sum $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}$ is well defined. By (6), we have
+$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\frac{1}{2}+i\int_{0}^{+\infty}\Big({\frac{1}{(1+it)^{s}}-\frac{1}{(1-it)^{s}}}\big)\frac{1}{e^{2\pi t}-1}\,\ dt,$
 $\frac{1}{2}+2\int_{0}^{+\infty}\frac{\sin(s\arctan(t))}{(1+t^{2})^{s/2}}\frac{1}{e^{2\pi t}-1}\ dt.$
@@
 $\zeta(s)=\sum_{m=0}^{+\infty}x_{m}P_{m}(s).$
-As we have seen in the Introduction the polynomials $P_{m}$ are related to the Laguerre polynomials by Mellin transform. We now give another approach to these polynomials.
 3.1 The sequence of polynomials $(P_{m})$
-We start with the observation that for $s\in\mathbb{C}$, and for every $n>0$, we have
-$\frac{1}{n^{s}}=\frac{1}{n+1}G_{s}(\frac{n-1}{n+1}),$ (18)
-where $G_{s}$ is the function defined on $D(0,1)=\{z\in\mathbb{C};|z|<1\}$ by
-$G_{s}:z\mapsto\frac{2}{1-z}\Big({\frac{1-z}{1+z}}\Big)^{s}$ (19)
-Since $\Re\big({\frac{1-z}{1+z}}\big)>0$ for $z\in D(0,1)$, we see that the function $G_{s}$ is analytic on $D(0,1)$. Let $P_{m}(s)$ denote the coefficients of its Taylor expansion
-$\frac{2}{1-z}\Big({\frac{1-z}{1+z}}\Big)^{s}=\sum_{m=0}^{+\infty}P_{m}(s)z^{m}\ \text{ for }\ \ \ z\in D(0,1)$ (20)
-Let $z=\frac{x}{1+x}$, with $x\in]0,1[$, then, using (20), we get
-$\sum_{m=0}^{+\infty}P_{m}(s)\Big(\frac{x}{1+x}\Big)^{m+1}=2x(1+2x)^{-s}=\sum_{k=0}^{+\infty}(-1)^{k}2^{k+1}\frac{(s)_{k}}{k!}x^{k+1},$ (21)
-where $(s)_{k}=\frac{\Gamma(s+k)}{\Gamma(s)}=s(s+1)\cdots(s+k-1)$ if $k\geq 1$ and $(s)_{0}=1.$ We use now the well-known equivalence related to Euler summation (cf.[H1])
+For $m\geq 0$}, let $\psi_{m}$ the function defined in the half-plane $\{Re(z)>0\}$ by
+$\psi_{m}(z)=\frac{1}{1+z}\big(\frac{z-1}{z+1}\big)^{m},$ (18)
+and let $(P_{m}(s))$ is the sequence of polynomials given by the generating function
+$\frac{2}{1-\xi}\Big({\frac{1-\xi}{1+\xi}}\Big)^{s}=\sum_{m=0}^{+\infty}P_{m}(s)\ \xi^{m}\ \text{ for }\ \ \ \xi\in D(0,1).$ (19)
+Then, with $\xi=\frac{z-1}{z+1}$, we have for $s\in\mathbb{C}$ and $Re(z)>0$, the expansion
+$z^{-s}=\sum_{m=0}^{+\infty}\psi_{m}(z)\,P_{m}(s).$ (20)
+To obtain an explicit expression of $P_{m}(s)$, let $\xi=\frac{x}{1+x}$, with $x\in]0,1[$, then by (19) we get
+$\sum_{m=0}^{+\infty}P_{m}(s)\Big(\frac{x}{1+x}\Big)^{m+1}=2x(1+2x)^{-s}=\sum_{k=0}^{+\infty}(-1)^{k}2^{k+1}\frac{(s)_{k}}{k!}x^{k+1},$
+where $(s)_{k}=\frac{\Gamma(s+k)}{\Gamma(s)}=s\cdots(s+k-1)$ if $k\geq 1$ and $(s)_{0}=1.$ Now the well-known equivalence (cf.[H1])
 $\sum_{m\geq 0}b_{m}(\frac{x}{1+x})^{m+1}=\sum_{k\geq 0}a_{k}x^{k+1}\ \Leftrightarrow\ b_{m}=\sum_{k=0}^{m}\binom{m}{k}\ a_{k},$
-which gives, for $P_{m}(s),$ the following explicit expression
+gives for $P_{m}(s)$ the following expression
 $P_{m}(s)=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k+1}\ \frac{(s)_{k}}{k!}.$ (21)
-We also have an integral expression for $P_{m}(s)$ when $0<\Re(s)<1$, since for $k\geq 0$, we can write
+It is easily verified that $P_{m}(s)$ is related, for $\Re(s)>0,$ with the Laguerre polynomial $L_{m}$ by
+$P_{m}(s)=\frac{2}{\Gamma(s)}\int_{0}^{+\infty}e^{-x}L_{m}(2x)x^{s-1}\ dx.$ (22)
+We have another integral expression for $P_{m}(s)$ when $0<\Re(s)<1$, since for $k\geq 0$, we can write
 $\frac{(s)_{k}}{k!}=\frac{1}{\Gamma(s)\Gamma(1-s)}\int_{0}^{1}t^{s-1+k}(1-t)^{-s}\ dt,$
@@
 $I_{a}\leq\int_{0}^{1}(1-u)^{a-1}u^{m}du+\int_{-1}^{0}(1+u)^{-a}(-u)^{m}\ du$ $=$ $\int_{0}^{1}(1-u)^{a-1}u^{m}du+\int_{0}^{1}(1-u)^{-a}u^{m}\ du.$
-Now if $0<\alpha<1$ we have, for every $x>0$
-$\int_{0}^{1}{(1-u)^{\alpha-1}}u^{x}du=\int_{0}^{+\infty}{e^{-t}}{(1-e^{-t})^{\alpha-1}}e^{-xt}dt\leq\int_{0}^{+\infty}t^{\alpha-1}e^{-xt}dt=\Gamma(\alpha)x^{-\alpha},$
-(the last inequality is justified by $1\leq\frac{t}{1-e^{-t}}\leq{e^{t}}$, for every $t>0$). Finally we obtain
+Now if $0<\alpha<1$ we have
+$\int_{0}^{1}{(1-u)^{\alpha-1}}u^{m}\ du=\int_{0}^{+\infty}{e^{-t}}{(1-e^{-t})^{\alpha-1}}e^{-mt}\ dt\leq\int_{0}^{+\infty}t^{\alpha-1}e^{-mt}\ dt=\Gamma(\alpha)\ m^{-\alpha},$
+(the last inequality is justified by $e^{-t}\leq\frac{1-e^{-t}}{t}\leq{1}$, for every $t>0$). With $\alpha=a$ and $\alpha=1-a$, we obtain
 $I_{a}\leq\frac{\Gamma(a)}{m^{a}}+\frac{\Gamma(1-a)}{m^{1-a}}.$
 Remark 6.
-It is easily verified (cf. [G-R]) that, for $m\geq 1$, we have $P_{m}(s)=2F(-m,s,1,2),$ where $F$ is the Gauss hypergeometric function
-$F(a,b,c,z)=\sum_{k\geq 0}\frac{(a)_{k}(b)_{k}}{(c)_{k}}\frac{z^{k}}{k!},$
+The polynomials $P_{m}$ are related with Gauss hypergeometric function
+$F(a,b,c,z)=\sum_{k\geq 0}\frac{(a)_{k}(b)_{k}}{(c)_{k}}\frac{z^{k}}{k!}.$
+It is easily verified with (21) that, for $m\geq 1$, we have
+$P_{m}(s)=2F(-m,s,1,2).$
 By classical properties of hypergeometric function (cf. [G-R]), we get the relations
@@
 $sP_{m}(s+1)-(s-1)P_{m}(s-1)=(2m+1)P_{m}(s).$ (25)
-3.2 Orthogonality relations
-Let $(\varphi_{m})_{m\geq 0}$ the sequence of functions
-$\varphi_{m}:x\mapsto\sqrt{2}\ e^{-x}L_{m}(2x).$
-It is well known (cf.
-[Leb]) that the sequence $(\varphi_{m})_{m\geq 0}$ is an orthonormal system in $(L^{2}(0,+\infty),dx)$.
-We verify as in the introduction that, by Mellin transform, we have for $\Re(s)>0$
-$\mathcal{M}\varphi_{m}(s)=\int_{0}^{+\infty}\varphi_{m}(x)x^{s-1}\ dx=\frac{1}{\sqrt{2}}{\Gamma(s)}P_{m}(s).$ (26)
-Thus the Mellin-Parseval formula (cf. [T]), gives
-$\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\mathcal{M}(\varphi_{m})(\frac{1}{2}+it)}\ \mathcal{M}(\varphi_{n})(\frac{1}{2}+it)\ dt=\int_{0}^{+\infty}\overline{\varphi_{m}(x)}\varphi_{n}(x)dx=\delta_{(m,n)},$
-with $\delta$ the usual Kronecker symbol. That is
-$\frac{1}{4\pi}\int_{-\infty}^{+\infty}\overline{\Gamma(\frac{1}{2}+it)}\overline{P_{m}(\frac{1}{2}+it)}\ \Gamma(\frac{1}{2}+it)P_{n}(\frac{1}{2}+it)dt=\delta_{(m,n)}.$
-Since
-$\Gamma(\frac{1}{2}+it)\overline{\Gamma(\frac{1}{2}+it)}=\frac{\pi}{\cosh\pi t},$
-then
-$\frac{1}{4}\int_{-\infty}^{+\infty}\overline{P_{m}(\frac{1}{2}+it)}\ P_{n}(\frac{1}{2}+it)\frac{dt}{\cosh\pi t}=\delta_{(m,n)}.$
-This proves that the sequence $(Q_{m})_{m\geq 0}$ of polynomials
-$Q_{m}:t\mapsto\frac{1}{2}P_{m}(\frac{1}{2}+it)=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k}\ \frac{(\frac{1}{2}+it)_{k}}{k!}$ (27)
-is an orthonormal system in $\big(L^{2}(-\infty,+\infty),\frac{dt}{\cosh(\pi t)}\big)$. It is also an orthonormal basis of this space since for $|a|<\pi$, we have
-$\int_{-\infty}^{+\infty}e^{a|t|}\frac{dt}{\cosh(\pi t)}<+\infty$
-By a classical result on orthogonal polynomials we deduce that, for every $m\geq 1$, the roots of $Q_{m}$ are real. This implies that the roots of $P_{m}$ are on the line $\{\Re(s)=\frac{1}{2}\}$.
-For the first values of $m$ we get
-$Q_{0}(t)$ $=$ $1,$
-$Q_{1}(t)$ $=$ $i(-2t),$
-$Q_{2}(t)$ $=$ $-2t^{2}+\frac{1}{2},$
-$Q_{3}(t)$ $=$ $i(\frac{4}{3}t^{3}-\frac{5}{3}t),$
-$Q_{4}(t)$ $=$ $\frac{2}{3}t^{4}-\frac{7}{3}t^{2}+\frac{3}{8}.$
-Remark 7.
-By (24) and (25) we have, for every $m\geq 0$, the relations
-$\overline{Q_{m}(t)}=Q_{m}(-t)=(-1)^{m}Q_{m}(t),$
-$\big(\frac{1}{2}+it\big)Q_{m}(t-i)+\big(\frac{1}{2}-it\big)Q_{m}(t+i)=(2m+1)Q_{m}(t).$
 3.2 Expansion of $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}$
-We now apply Ramanujan’s summation on both sides of (18). Using (20), we get
-$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}G_{s}(\frac{n-1}{n+1})=\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s).$
-The following theorem says that, for $0<\Re(s)<1$, these last two sums can be inverted.
 Theorem 6.
-For $0<\Re(s)<1$ we have
-$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\sum_{m=0}^{+\infty}z_{m}P_{m}(s)\ \text{ where }\ z_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}.$
+For $0<\Re(s)<1,$ we have
+$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\sum_{m=0}^{+\infty}z_{m}P_{m}(s)\ \text{ where }\ z_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}.$ (26)
 And we obtain the expansion
@@
 Proof.
-By definition 2, we can write
+We apply (20) with $z=n$ integer $\geq 1$, this gives
+$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}\psi_{m}(n)P_{m}(s)=\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s).$ (28)
+We now prove that, for $0<\Re(s)<1$, these last two sums can be inverted, that is to say, the series
+$\sum_{m\geq 0}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s)$ (29)
+is convergent for every $0<\Re(s)<1$, and its sum is equal to
+$\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s).$
+By definition (6), we can write
 $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}=c_{m}+i\int_{0}^{+\infty}\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\frac{1}{e^{2\pi t}-1}\ dt$
-with $c_{0}=1/4$ if $m=0$ and $c_{m}=0$ if $m\geq 1$. We will now prove that the series
-$\sigma(s)=\sum_{m\geq 0}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s)$
-is convergent for every $0<\Re(s)<1$, and that its sum is equal to
-$\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s).$
-Since $c_{0}P_{0}(s)=1/2$, then, by definition 2, the series $\sigma(s)$ is also
-$\sigma(s)=\frac{1}{2}+i\sum_{m\geq 0}\int_{0}^{+\infty}\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\frac{1}{e^{2\pi t}-1}\ dt\ P_{m}(s).$ (29)
+with $c_{0}=1/4$ if $m=0$ and $c_{m}=0$ if $m\geq 1$. Since $c_{0}P_{0}(s)=1/2$, then the series (29) is
+$\frac{1}{2}+i\sum_{m\geq 0}\int_{0}^{+\infty}\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\frac{1}{e^{2\pi t}-1}\ dt\ P_{m}(s).$
 To prove the convergence of this series and invert signs $\sum$ and $\int$ we use the following lemma.
@@
 $\sum_{m\geq 0}\Big|\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\ P_{m}(s)\Big|$ (30)
-is convergent and
-$\int_{0}^{+\infty}\sum_{m=0}^{+\infty}\Big|\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\ P_{m}(s)\Big|\ \frac{1}{e^{2\pi t}-1}dt<+\infty.$ (31)
-Proof of the lemma.
+is convergent for every $t>0$, and its sum verify
+$\int_{0}^{+\infty}\sum_{m=0}^{+\infty}\Big|\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\ P_{m}(s)\Big|\ \frac{1}{e^{2\pi t}-1}\ dt<+\infty.$
+By this lemma the series (29) is convergent for every $0<\Re(s)<1$, and we get
+$\sum_{m=0}^{+\infty}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s)$
+$\frac{1}{2}+i\int_{0}^{+\infty}\Big(\frac{1}{2+it}\sum_{m=0}^{+\infty}\big(\frac{it}{2+it}\big)^{m}P_{m}(s)-\frac{1}{2-it}\sum_{m=0}^{+\infty}\big(\frac{-it}{2-it}\big)^{m}P_{m}(s)\Big)\frac{1}{e^{2\pi t}-1}\ dt.$
+Using (19) with $\xi=\frac{\pm it}{2\pm it}$, this last expression is equal to
+$\frac{1}{2}+i\int_{0}^{+\infty}\Big(\big(\frac{1}{1+it}\big)^{s}-\big(\frac{1}{1-it}\big)^{s}\Big)\frac{1}{e^{2\pi t}-1}\ dt=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}.$
+The relation (27) follows immediately from (17).
+Proof of lemma 2.
 We have
@@
 $\frac{1}{\sqrt{t^{2}+4}}\sum_{m=2}^{+\infty}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}\frac{1}{m^{\alpha}}\leq\frac{1}{\sqrt{t^{2}+4}}\sum_{m=1}^{+\infty}\big(\frac{t}{\sqrt{t^{2}+4}}\big)^{m}\leq\frac{t}{{t^{2}+4}}\Big(\frac{1}{1-\frac{t}{\sqrt{t^{2}+4}}}\Big)\leq\frac{t^{2}}{4}.$
-End of the proof of the lemma.
-By the preceding lemma the series $\sigma(s)$ is convergent for every $0<\Re(s)<1$, and we get
-$\sum_{m=0}^{+\infty}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s)$
-$\frac{1}{2}+i\int_{0}^{+\infty}\Big(\frac{1}{2+it}\sum_{m=0}^{+\infty}\big(\frac{it}{2+it}\big)^{m}P_{m}(s)-\frac{1}{2-it}\sum_{m=0}^{+\infty}\big(\frac{-it}{2-it}\big)^{m}P_{m}(s)\Big)\frac{1}{e^{2\pi t}-1}\ dt.$
-Using (20) with $z=\frac{\pm it}{2\pm it}$, this last expression is equal to
-$\frac{1}{2}+i\int_{0}^{+\infty}\Big(\big(\frac{1}{1+it}\big)^{s}-\big(\frac{1}{1-it}\big)^{s}\Big)\frac{1}{e^{2\pi t}-1}\ dt,$
-which is by definition 2, also equal to
-$\frac{1}{2}+\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}-\frac{1}{2}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}.$
-The relation (28) follows immediately from (17).
-3.4 Evaluation of the sum $z_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}$
-If we apply the shift formula of corollary 1 to the function $f:x\mapsto\frac{1}{x}\big(1-\frac{2}{x}\big)^{m}$, we get
-$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}\Big(1-\frac{2}{n}\Big)^{m}-(-1)^{m}+y_{m}$
-with
-$y_{m}=\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}dx=\int_{1}^{2}\Big(1-\frac{2}{x}\Big)^{m}\frac{1}{x}dx.$
-Now the binomial expansion
+3.3 Evaluation of the sum $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}$
+Lemma 3.
+For any integer $m\geq 0$
+$z_{m}:=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}=x_{m}+y_{m}$
+with $x_{0}=\gamma-1,\ \ y_{0}=\log(2),$ and for $m\geq 1$
+$x_{m}=\gamma+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\Big(\zeta(k+1)-\frac{1}{k}\Big)-(-1)^{m},$ (31)
+and
+$y_{m}=\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}\ \ dx=\log(2)+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\frac{1}{k}\Big(1-\frac{1}{2^{k}}\Big).$ (32)
+Proof.
+We apply the shift formula of corollary 1 to the function $f:x\mapsto\frac{1}{x}\big(1-\frac{2}{x}\big)^{m}$, then we get
+$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}\Big(1-\frac{2}{n}\Big)^{m}-(-1)^{m}+\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}\ dx$
+For any $m$ integer $\geq 0$, let
+$x_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}\Big(1-\frac{2}{n}\Big)^{m}-(-1)^{m}\ \text{ and }\ y_{m}=\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}\ dx$
+We have $x_{0}=\gamma-1,\ \ y_{0}=\log(2).$ And by the binomial expansion
 $\Big(1-\frac{2}{x}\Big)^{m}=\sum_{k=0}^{m}\binom{m}{k}(-2)^{k}\frac{1}{x^{k}}$
-gives
-$z_{m}=\sum_{k=0}^{m}\binom{m}{k}(-2)^{k}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{k+1}}-(-1)^{m}+\sum_{k=0}^{m}\binom{m}{k}(-2)^{k}\int_{1}^{2}\frac{1}{x^{k+1}}dx.$
-Thus we obtain $z_{m}=x_{m}+y_{m}$, where $x_{0}=\gamma-1,\ \ y_{0}=\log(2),$ and for $m\geq 1$
-$x_{m}$ $=$ $\gamma+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\Big(\zeta(k+1)-\frac{1}{k}\Big)-(-1)^{m},$ (32)
-$y_{m}$ $=$ $\log(2)+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\frac{1}{k}\Big(1-\frac{1}{2^{k}}\Big).$
-For the first values of $m$ we get
-$x_{0}$ $=$ $\gamma-1,$
-$x_{1}$ $=$ $\gamma-2\zeta(2)+3,$
-$x_{2}$ $=$ $\gamma-4\zeta(2)+4\zeta(3)+1,$
-$x_{3}$ $=$ $\gamma-6\zeta(2)+12\zeta(3)-8\zeta(4)+11/3,$
-$x_{4}$ $=$ $\gamma-8\zeta(2)+24\zeta(3)-32\zeta(4)+16\zeta(5)+5/3.$
+we get for $m\geq 1$
+$x_{m}$ $=$ $\gamma+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{k}}-(-1)^{m}.$
+and
+$y_{m}=\int_{1}^{2}\frac{1}{x}\big(1-\frac{2}{x}\big)^{m}\ \ dx=\sum_{k=0}^{m}\binom{m}{k}(-2)^{k}\int_{1}^{2}\frac{1}{x^{k+1}}\ dx.$
+The conclusion follows from (17) which gives for $k\geq 1$
+$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{k}}=\zeta(k+1)-\frac{1}{k}.$
+3.4 An expansion of Zeta
 Theorem 7.
@@
 Proof.
-This follows from the expansion (28) of theorem 6, and the combinatorial identity
+This follows from the expansion (27) of theorem 6. Since by lemma 3, we have
+$z_{m}=x_{m}+y_{m}=\gamma+\log(2)+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\zeta(k+1)+(-1)^{m+1}+\sum_{k=1}^{m}\binom{m}{k}\frac{(-1)^{k-1}}{k},$
+and we can use the combinatorial identity
 $\sum_{k=1}^{m}\binom{m}{k}\frac{(-1)^{k-1}}{k}=\sum_{k=1}^{m}\frac{1}{k}.$
-3.5 An expansion of Zeta
 The previous theorem will enable us to get an expansion of Zeta in the strip $\{0<\Re(s)<1\}$. Indeed, the polar part 1/(s1)1/(s-1) of (33) has the simple expansion given by the following lemma.
@@
 For $0<\Re(s)<1$, we write
-$\sum_{m=0}^{+\infty}y_{m}P_{m}(s)=\sum_{m=0}^{+\infty}\Big(\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}dx\Big)P_{m}(s)=\int_{0}^{1}\sum_{m=0}^{+\infty}\Big(\frac{x-1}{x+1}\Big)^{m}P_{m}(s)\frac{1}{x+1}dx.$ (35)
+$\sum_{m=0}^{+\infty}y_{m}P_{m}(s)=\sum_{m=0}^{+\infty}\Big(\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}\ \ dx\Big)P_{m}(s)=\int_{0}^{1}\sum_{m=0}^{+\infty}\frac{1}{x+1}\Big(\frac{x-1}{x+1}\Big)^{m}P_{m}(s)\ dx.$ (35)
 This interchanging of signs $\sum$ and $\int$ is justified by
 $\sum_{m=0}^{+\infty}\int_{0}^{1}|\frac{x-1}{x+1}|^{m}\ |P_{m}(s)|\frac{1}{x+1}\ dx\leq\sum_{m=0}^{+\infty}\int_{0}^{1}\frac{1}{(x+1)^{m+1}}\ dx\ |P_{m}(s)|\leq\sum_{m=0}^{+\infty}\frac{1}{m}|P_{m}(s)|,$
-and we immediately check, with the lemma 1, that the latter series is convergent.
-The relations (18) and (20) allows us to write for every $x>0$
-$\sum_{m=0}^{+\infty}\frac{1}{x+1}\Big(\frac{x-1}{x+1}\Big)^{m}P_{m}(s)=\frac{1}{x^{s}},$
-and by (35) we have
-$\sum_{m=0}^{+\infty}y_{m}P_{m}(s)=\int_{0}^{1}\frac{1}{x^{s}}dx=-\frac{1}{s-1}.$
+and we immediately check, with lemma 1, that the latter series is convergent.
+The relation (34) is then a consequence of (20) and (35).
 Theorem 8.
@@
 $\zeta(s)=\sum_{m=0}^{+\infty}x_{m}\,P_{m}(s)$ (36)
-with
+with $x_{0}=\gamma-1,$ and for $m\geq 1$
 $x_{m}=\gamma+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\zeta(k+1)+r_{m}$ (37)
 and $r_{m}$ are the rational numbers
-$r_{m}=(-1)^{m+1}+\sum_{k=1}^{m}\frac{1+(-1)^{k-1}}{k}$
+$r_{m}=(-1)^{m+1}+\sum_{k=1}^{m}\frac{1+(-1)^{k-1}}{k}.$
 Proof.
@@
 $\sum_{k=1}^{m}\binom{m}{k}\frac{(-2)^{k-1}}{k}=\sum_{k=1}^{m}\frac{1+(-1)^{k-1}}{k}.$
+For the first values of $m$ we get
+$x_{1}$ $=$ $\gamma-2\zeta(2)+3,$
+$x_{2}$ $=$ $\gamma-4\zeta(2)+4\zeta(3)+1,$
+$x_{3}$ $=$ $\gamma-6\zeta(2)+12\zeta(3)-8\zeta(4)+11/3$
+$x_{4}$ $=$ $\gamma-8\zeta(2)+24\zeta(3)-32\zeta(4)+16\zeta(5)+5/3.$
+3.5 Orthogonality relations
+The polynomial $P_{m}$ is related to the Mellin transform of the function
+$\varphi_{m}:x\mapsto\sqrt{2}\ e^{-x}L_{m}(2x),$ (38)
+since we verify easily with (22) that, for $\Re(s)>0$, we have
+$\mathcal{M}\varphi_{m}(s)=\int_{0}^{+\infty}\varphi_{m}(x)x^{s-1}\ \ dx=\frac{1}{\sqrt{2}}{\Gamma(s)}P_{m}(s).$ (39)
+It is well known (cf. [Leb]) that the sequence $(\varphi_{m})_{m\geq 0}$ is an orthonormal system in $(L^{2}(0,+\infty),\ dx)$.
+Theorem 9.
+The sequence $(Q_{m})_{m\geq 0}$ of polynomials
+$Q_{m}:t\mapsto\frac{1}{2}P_{m}(\frac{1}{2}+it)$ (40)
+is an orthonormal system in $\big(L^{2}(\mathbb{R}),\frac{dt}{\cosh(\pi t)}\big)$. This implies that the roots of $P_{m}$ are on the line $\{\Re(s)=\frac{1}{2}\}$.
+Proof.
+By the Mellin-Parseval formula (cf. [T]), we have
+$\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\mathcal{M}(\varphi_{m})(\frac{1}{2}+it)}\ \mathcal{M}(\varphi_{n})(\frac{1}{2}+it)\ dt=\int_{0}^{+\infty}\overline{\varphi_{m}(x)}\varphi_{n}(x)\ dx=\delta_{(m,n)},$
+with $\delta$ the usual Kronecker symbol. That is
+$\frac{1}{4\pi}\int_{-\infty}^{+\infty}\overline{\Gamma(\frac{1}{2}+it)}\overline{P_{m}(\frac{1}{2}+it)}\ \Gamma(\frac{1}{2}+it)P_{n}(\frac{1}{2}+it)\ dt=\delta_{(m,n)}.$
+Since
+$\Gamma(\frac{1}{2}+it)\overline{\Gamma(\frac{1}{2}+it)}=\frac{\pi}{\cosh\pi t},$
+then
+$\frac{1}{4}\int_{-\infty}^{+\infty}\overline{P_{m}(\frac{1}{2}+it)}\ P_{n}(\frac{1}{2}+it)\frac{dt}{\cosh\pi t}=\delta_{(m,n)}.$
+By a classical result on orthogonal polynomials we deduce that, for every $m\geq 1$, the roots of $Q_{m}$ are real. Thus the roots of $P_{m}$ are on the line $\{\Re(s)=\frac{1}{2}\}$. ∎
+Remark 7.
+By (24) we have, for every $m\geq 0$, the relations
+$\overline{Q_{m}(t)}=Q_{m}(-t)=(-1)^{m}Q_{m}(t),$
+thus we have for $t\in\mathbb{R}$
+$\Re\zeta(\frac{1}{2}+it)$ $=$ $2\sum_{m=0}^{+\infty}x_{2m}Q_{2m}(t)$
+$\Im\zeta(\frac{1}{2}+it)$ $=$ $2\sum_{m=0}^{+\infty}x_{2m+1}Q_{2m+1}(t).$
+By (25) we have,
+$TQ_{m}=(2m+1)Q_{m},$
+where $T$ is the operator on $\mathbb{R}[t]$ defined by
+$Tf(t)=\big(\frac{1}{2}+it\big)f(t-i)+\big(\frac{1}{2}-it\big)f(t+i).$
 3.6 An orthonormal expansion
-Let $X(0)=\gamma-1$ and for $k\geq 1$
+Let $X(0)=\gamma-1$ and for $k\geq 1$ let
 $X(k)=\zeta(k+1)-1-\frac{1}{k}.$
@@
 $x_{m}=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k}X(k).$
-Let
+Let the function $f_{0}$ defined for $t>0$ by
 $f_{0}(t)=\frac{1}{e^{t}-1}-\frac{1}{t},$
-then we have for $k\geq 0$
-$X(k)=\int_{0}^{+\infty}e^{-t}\frac{t^{k}}{k!}\ f_{0}(t)dt$
-By definition of Laguerre polynomials, we obtain for the coefficients $x_{m}$ the integral expression
+it is well-known that $\mathcal{M}f_{0}(s)=\Gamma(s)\zeta(s)$ for $0<\Re(s)<1$. We easily verify that for $k\geq 0$ we have
+$X(k)=\int_{0}^{+\infty}e^{-t}\frac{t^{k}}{k!}\ f_{0}(t)\ dt.$
+Then by definition of Laguerre polynomials, we obtain
 $x_{m}=\int_{0}^{+\infty}e^{-t}L_{m}(2t)f_{0}(t)\ dt=\frac{1}{\sqrt{2}}\int_{0}^{+\infty}\varphi_{m}(t)f_{0}(t)\ dt=\frac{1}{\sqrt{2}}(\varphi_{m},f_{0}).$ (41)
+This means that we have, in $L^{2}(]0,+\infty[)$, the orthonormal expansion
+$f_{0}=\sqrt{2}\,\sum_{m=0}^{+\infty}x_{m}\,\varphi_{m}.$ (42)
 Theorem 10.
-We have for any integer $m\geq 0$
+For any integer $m\geq 0$, we have
 $x_{m}=\frac{1}{2}\int_{-\infty}^{+\infty}\overline{Q_{m}(t)}\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}.$ (43)
 Proof.
-By Mellin-Parseval formula (cf. [T]), we have
-$\frac{1}{\sqrt{2}}\int_{0}^{+\infty}{\varphi_{m}(t)}f_{0}(t)dt=\frac{1}{2\pi\sqrt{2}}\int_{-\infty}^{+\infty}\overline{\mathcal{M}(\varphi_{m})(\frac{1}{2}+it)}\ \mathcal{M}(f_{0})(\frac{1}{2}+it)\ dt$
-Since (cf. [T]) we have $\mathcal{M}f(s)=\Gamma(s)\zeta(s)$ for $0<\Re(s)<1$, then, using (26), we get
-$x_{m}=\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\Gamma(\frac{1}{2}+it)\frac{1}{2}P_{m}(\frac{1}{2}+it)}\Gamma((\frac{1}{2}+it))\frac{1}{2}\zeta(\frac{1}{2}+it)\ dt,$
-since $Q_{m}(t)=\frac{1}{2}P_{m}(\frac{1}{2}+it)$, this is the integral expression (39).
+By (42) and Mellin-Parseval formula (cf. [T]), we have
+$x_{m}=\frac{1}{\sqrt{2}}\int_{0}^{+\infty}{\varphi_{m}(t)}f_{0}(t)\ dt=\frac{1}{2\pi\sqrt{2}}\int_{-\infty}^{+\infty}\overline{\mathcal{M}(\varphi_{m})(\frac{1}{2}+it)}\ \mathcal{M}(f_{0})(\frac{1}{2}+it)\ dt,$
+thus we get
+$x_{m}$ $=$ $\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\Gamma(\frac{1}{2}+it)\frac{1}{2}P_{m}(\frac{1}{2}+it)}\Gamma((\frac{1}{2}+it))\frac{1}{2}\zeta(\frac{1}{2}+it)\ dt$
+$=$ $\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\frac{1}{2}P_{m}(\frac{1}{2}+it)}\frac{1}{2}\zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}.$
+Remark 8.
 The previous theorem means that the pointwise convergent expansion
 $\zeta(\frac{1}{2}+it)=2\sum_{m=0}^{+\infty}x_{m}\,Q_{m}(t)$
-is also an orthonormal expansion in the space $\big(L^{2}(-\infty,+\infty),\frac{dt}{\cosh(\pi t)}\big)$.
-Then we have
+is also an orthonormal expansion in the space $L^{2}\big(\mathbb{R},\frac{dt}{\cosh(\pi t)}\big)$. Then we have
 $\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=4\sum_{m=0}^{+\infty}x_{m}^{2}.$ (44)
-We give next a formal evaluation of this last sum using Ramanujan’s summation.
-3.7 A generating function
-Let $\mathcal{L}$ the Laplace transform
-$\mathcal{L}f(z)=\int_{0}^{+\infty}e^{-xz}f(x)\ dx=(e^{-xz},f).$
-By Paley-Wiener theorem (cf.
-[Ho], [P]) , the transformation $\frac{1}{\sqrt{2\pi}}\mathcal{L}$ maps isometrically the space $L^{2}(]0,+\infty[,dx)$ to the Hardy space $H^{2}(\mathbb{C}_{+})$ of analytic functions $g$ in the half-plane $\mathbb{C}_{+}=\{z|\Re(z>0\}$, such that : there exists $M_{g}>0$ with
-$\int_{\mathbb{R}}|g(x+iy)|^{2}\ dy\leq M_{g}\text{ for all }x>0.$
-Each such $g$ has non-tangential limits $g(iy)=\lim_{x\to 0}g(x+iy)$ at almost every point of the imaginary axis. The space $H^{2}(\mathbb{C}_{+})$ is an Hilbert space with the inner product
-$(g,h)_{H^{2}}=\int_{\mathbb{R}}\overline{g(iy)}{h(iy)}dy,$
-and the functions $\psi_{m}$ $(m\geq 0)$, form an orthonormal basis of $H^{2}(\mathbb{C}_{+})$. If $f\in L^{2}(]0,+\infty[,dx)$. From
-$e^{-xz}=\sqrt{2}\ \sum_{k=0}^{+\infty}\varphi_{k}(x)\psi_{k}(z).$
-we get
-$\frac{1}{\sqrt{2\pi}}\mathcal{L}\varphi_{m}(z)=\frac{1}{\sqrt{2\pi}}(e^{-xz},\varphi_{m})=\frac{1}{\sqrt{\pi}}\psi_{m}(z)$
-The system $(\frac{1}{\sqrt{\pi}}\psi_{m})_{m\geq 0}$ is an orthonormal basis of $H^{2}(\mathbb{C}_{+})$. From the expansion
-$f_{0}=\sum_{m\geq 0}(\varphi_{m},f_{0})\varphi_{m}$
-we get the expansion
-$\frac{1}{\sqrt{2\pi}}\mathcal{L}f_{0}(z)=\sum_{m\geq 0}(\varphi_{m},f_{0})\frac{1}{\sqrt{\pi}}\psi_{m}(z).$
-and, by (38) we deduce
-$\mathcal{L}f_{0}(z)=2\sum_{m\geq 0}x_{m}\psi_{m}(z).$
-Since it is well known (cf. [G-R]) that
-$\int_{0}^{+\infty}e^{-xz}f_{0}(x)dx=\log(z)-\Psi(z)-\frac{1}{z}$
-then we get
+We now give a formal evaluation of this last sum using Ramanujan’s summation. Let $\mathcal{L}$ the Laplace transform
+$\mathcal{L}f(z)=\int_{0}^{+\infty}e^{-xz}f(x)\ dx.$
+We have (cf. [G-R]) for $Re(z)>0$, the Laplace transforms
+$\mathcal{L}\varphi_{m}(z)$ $=$ ${\sqrt{2}}\ \psi_{m}(z),$
+$\mathcal{L}f_{0}(z)$ $=$ $\log(z)-\Psi(z)-\frac{1}{z}.$
+By Laplace transform of expansion (42) we get
 $2\sum_{m=0}^{+\infty}x_{m}\psi_{m}(z)=\log(z)-\Psi(z)-\frac{1}{z}.$ (45)
-Equivalently this gives the generating function
-$\sum_{m=0}^{+\infty}x_{m}u^{m}=\frac{1}{1-u}\Big(\log(\frac{1+u}{1-u})-\Psi(\frac{1+u}{1-u})\Big)-\frac{1}{1+u}.$ (42)
-Remark 8.
 Now, if we apply (45) with $z=1,2,3,\dots$ and take Ramanujan’s summation, we get
 $2\sum_{m=0}^{+\infty}x_{m}\sum_{n\geq 1}^{\mathcal{R}}\psi_{m}(n)=2\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}x_{m}\psi_{m}(n)=\sum_{n\geq 1}^{\mathcal{R}}\log(n)-\sum_{n\geq 1}^{\mathcal{R}}\Psi(n)-\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}.$
-Then, with the results of examples 4-6, this relation gives
+Then, with the evaluations of these sums given in lemma 3 and in (13), (14), (15), we get
 $2\sum_{m=0}^{+\infty}x_{m}(x_{m}+y_{n})=\log(2\pi)-\gamma-\frac{3}{2}.$
-Since $y_{m}=\int_{0}^{1}\psi_{m}(x)dx$ then by (41) we have
+Since by (32) $y_{m}=\int_{0}^{1}\psi_{m}(x)\ dx$ then using (45) we have
 $2\sum_{m=0}^{+\infty}x_{m}y_{n}=\int_{0}^{1}2\sum_{m=0}^{+\infty}x_{m}\psi_{m}(x)\ dx=\int_{0}^{1}(\log(x)-\Psi(1+x))\ dx=-1.$
@@
 $\sum_{m=0}^{+\infty}x_{m}^{2}=\frac{1}{2}\log(2\pi)-\frac{1}{2}\gamma-\frac{1}{4}.$
-Then (40) gives us the relation
-$\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=2\log(2\pi)-2\gamma-1,$ (43)
+With (42) and (44) this gives
+$\int_{0}^{+\infty}\Big(\frac{1}{e^{t}-1}-\frac{1}{t}\Big)^{2}\ dt=\frac{1}{2}\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=\log(2\pi)-\gamma-\frac{1}{2},$ (46)

### v3

@@
 To avoid functions like $z\mapsto\sin(\pi z)$, we will consider a space of analytic functions such that Ramanujan’s constant, defined by (4), depends only on the values $f(n)$ for all integers $n\geq 1$}.
-2.3 The space $\mathcal{E}^{\pi}$
+2.3 The space $\mathcal{E}$
 Definition 1.
@@
 Theorem 3.
-Let $f\in\mathcal{E}^{\pi}$. There is a unique function $\varphi_{f}\in\mathcal{E}^{\pi}$, such that
+Let $f\in\mathcal{E}$. There is a unique function $\varphi_{f}\in\mathcal{E}$, such that
 […]
@@
 Then formula (31) can be written as a binomial transform
-$x_{m}=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k}X(k).$
+$x_{m}=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k}X(k).$ (41)
 Let the function $f_{0}$ defined for $t>0$ by
@@
 Then by definition of Laguerre polynomials, we obtain
-$x_{m}=\int_{0}^{+\infty}e^{-t}L_{m}(2t)f_{0}(t)\ dt=\frac{1}{\sqrt{2}}\int_{0}^{+\infty}\varphi_{m}(t)f_{0}(t)\ dt=\frac{1}{\sqrt{2}}(\varphi_{m},f_{0}).$ (41)
+$x_{m}=\int_{0}^{+\infty}e^{-t}L_{m}(2t)f_{0}(t)\ dt=\frac{1}{\sqrt{2}}\int_{0}^{+\infty}\varphi_{m}(t)f_{0}(t)\ dt=\frac{1}{\sqrt{2}}(\varphi_{m},f_{0}).$
 This means that we have, in $L^{2}(]0,+\infty[)$, the orthonormal expansion
@@
 thus we get
-$x_{m}$ $=$ $\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\Gamma(\frac{1}{2}+it)\frac{1}{2}P_{m}(\frac{1}{2}+it)}\Gamma((\frac{1}{2}+it))\frac{1}{2}\zeta(\frac{1}{2}+it)\ dt$
-$=$ $\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\frac{1}{2}P_{m}(\frac{1}{2}+it)}\frac{1}{2}\zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}.$
+$x_{m}$ $=$ $\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\Gamma(\frac{1}{2}+it)\frac{1}{2}P_{m}(\frac{1}{2}+it)}\ \Gamma((\frac{1}{2}+it))\zeta(\frac{1}{2}+it)\ dt$
+$=$ $\frac{1}{2}\int_{-\infty}^{+\infty}\overline{\frac{1}{2}P_{m}(\frac{1}{2}+it)}\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}.$
+Corollary 2.
+We have
+$\gamma=1+\frac{1}{2}\int_{-\infty}^{+\infty}\zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)},$
+and for $m\geq 1$
+$\zeta(m+1)=1+\frac{1}{m}+\frac{1}{2}\int_{-\infty}^{+\infty}\zeta(\frac{1}{2}+it)\frac{(\frac{1}{2}-it)_{m}}{m!}\frac{dt}{\cosh(\pi t)}.$
+Proof.
+With (41) and inversion of the binomial transform, we get
+$\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}x_{k}=2^{m}X(m),$
+and from (21)
+$\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}\overline{Q_{k}(t)}=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}\overline{\frac{1}{2}P_{k}(\frac{1}{2}+it)}=2^{m}\ \frac{(\frac{1}{2}-it)_{m}}{m!}.$
+Thus by (43) we obtain the integral expression
+$X(m)$ $=$ $\frac{1}{2^{m+1}}\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}\int_{-\infty}^{+\infty}\overline{Q_{k}(t)}\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}$
+$=$ $\frac{1}{2^{m+1}}\int_{-\infty}^{+\infty}\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}\overline{Q_{k}(t)}\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}$
+$=$ $\frac{1}{2}\int_{-\infty}^{+\infty}\frac{(\frac{1}{2}-it)_{m}}{m!}\zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}.$
+Example 8.
+For $m=1,2,$ we have
+$\zeta(2)$ $=$ $2+\frac{1}{2}\int_{-\infty}^{+\infty}(\frac{1}{2}-it)\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}$
+$\zeta(3)$ $=$ $\frac{3}{2}-\frac{1}{4}\int_{-\infty}^{+\infty}(t^{2}+2it-\frac{3}{4})\ \zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}.$
 Remark 8.
@@
 Then, with the evaluations of these sums given in lemma 3 and in (13), (14), (15), we get
-$2\sum_{m=0}^{+\infty}x_{m}(x_{m}+y_{n})=\log(2\pi)-\gamma-\frac{3}{2}.$
+$2\sum_{m=0}^{+\infty}x_{m}(x_{m}+y_{m})=\log(2\pi)-\gamma-\frac{3}{2}.$
 Since by (32) $y_{m}=\int_{0}^{1}\psi_{m}(x)\ dx$ then using (45) we have
-$2\sum_{m=0}^{+\infty}x_{m}y_{n}=\int_{0}^{1}2\sum_{m=0}^{+\infty}x_{m}\psi_{m}(x)\ dx=\int_{0}^{1}(\log(x)-\Psi(1+x))\ dx=-1.$
+$2\sum_{m=0}^{+\infty}x_{m}y_{m}=\int_{0}^{1}2\sum_{m=0}^{+\infty}x_{m}\psi_{m}(x)\ dx=\int_{0}^{1}(\log(x)-\Psi(1+x))\ dx=-1.$
 Finally we get
 $\sum_{m=0}^{+\infty}x_{m}^{2}=\frac{1}{2}\log(2\pi)-\frac{1}{2}\gamma-\frac{1}{4}.$
-With (42) and (44) this gives
-$\int_{0}^{+\infty}\Big(\frac{1}{e^{t}-1}-\frac{1}{t}\Big)^{2}\ dt=\frac{1}{2}\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=\log(2\pi)-\gamma-\frac{1}{2},$ (46)
+With (44) this gives
+$\frac{1}{2}\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=\log(2\pi)-\gamma-\frac{1}{2}.$ (46)

### v4

@@
-A new expansion of the Riemann zeta function
+arXiv:2512.11405v4[math.NT] 01 Mar 2026
 Abstract
-We introduce a family of orthogonal polynomials $P_{m}(s)$, whose roots lie on the line $\{\Re(s)=1/2\}$, and we give an expansion of the Riemann zeta function in the critical strip as a convergent series $\sum_{m\geq 0}x_{m}P_{m}(s)$.
+We introduce a family of orthogonal polynomials $P_{m}(s)$, whose roots lie on the line $\{\Re(s)=1/2\}$, and we give an expansion of the Riemann’s zeta function in the critical strip as a convergent series $\sum_{m\geq 0}x_{m}P_{m}(s)$.
 The coefficients $x_{m}$of this expansion are expressed as linear combinations with rational coefficients of the Euler constant $\gamma$ and the values $\zeta(2),$ $\zeta(3),$ $\dots,$ $\zeta(m+1).$
 1 Introduction
-Some polynomial expansions of the Riemann zeta function (cf.[E-K-K], [F-V],[B],[M], [V]) are based on the Newton interpolation formula for functions related to Zeta. For example
+Some polynomial expansions of the Riemann’s zeta function (cf.[E-K-K], [F-V],[B],[M], [V]) are based on the Newton interpolation formula for functions related to the zeta function. For example
 $\zeta(s)-\frac{1}{s-1}=\sum_{m=0}^{+\infty}(-1)^{m}c_{m}\binom{s-1}{m}$,
@@
 where $\gamma_{m}$ are the Stieltjes constants.
-The polynomial expansion of the Riemann zeta function that we give in this paper is related to the Mellin transform of the functions $x\mapsto e^{-x}L_{m}(2x)$ where $(L_{m})_{m\geq 0}$ is the sequence of Laguerre polynomials defined by
+The polynomial expansion of the Riemann’s zeta function that we give in this paper is related to the Mellin transform of the functions $x\mapsto e^{-x}L_{m}(2x)$ where $(L_{m})_{m\geq 0}$ is the sequence of Laguerre polynomials defined by
 $L_{m}(x)=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}\frac{x^{k}}{k!}.$
@@
 $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\zeta(s)-\frac{1}{s-1},$
-then, after an explicit evaluation of $z_{n}$, we obtain, for $0<\Re(s)<1$, the following expansion
-$\zeta(s)=\sum_{m=0}^{+\infty}x_{m}P_{m}(s)$
+then, after an explicit evaluation of $z_{m}$, we obtain, for $0<\Re(s)<1$, the following expansion
+$\zeta(s)=\sum_{m=0}^{+\infty}x_{m}P_{m}(s),$
 with
 $x_{m}=\gamma+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\zeta(k+1)+r_{m}$
-and $r_{m}$ are the rational numbers
+where $r_{m}$ is given by
 $r_{m}=(-1)^{m+1}+\sum_{k=1}^{m}\frac{1+(-1)^{k-1}}{k}.$
@@
 b(x)=x[x]12 and [x] denote the integer part of x.b(x)=x-[x]-$\frac${1}{2}\ $\text${ and }\ [x]\ $\text${ denote the integer part of }\ x.
-Let’s assume that the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\ dx$ is convergent. The previous formula gives
+Let’s assume that the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\ dx$ is convergent. Then the previous formula gives
 $\sum_{k=1}^{n}f(k)=\int_{1}^{n}f(x)\,dx+C(f)+\frac{1}{2}f(n)-\int_{n}^{+\infty}b(x)f^{\prime}(x)\ dx$}, (2)
  $\text{ with }\ \ C(f)=\frac{1}{2}f(1)+\int_{1}^{+\infty}b(x)f^{\prime}(x)\ dx.$
-The constant $C(f)$ is called Ramanujan’s constant of the series $\sum_{n\geq 1}f(n)$ (cf. [H1]).
+The constant $C(f)$ is called the Ramanujan constant of the series $\sum_{n\geq 1}f(n)$ (cf. [H1]).
 Clearly $C(f)$ depends linearly of $f$. And if the series $\sum_{n\geq 1}f(n)$ is convergent, then by (2) we have
@@
 $\sum_{k=1}^{n}\frac{1}{k}=\log n+C(f)+\frac{1}{2n}+\int_{n}^{+\infty}b(x)\frac{1}{x^{2}}\ dx,$
-and we see that the Ramanujan’s constant of the series $\sum_{n\geq 1}\frac{1}{n}$ is the Euler constant :
+and we see that the Ramanujan constant of the series $\sum_{n\geq 1}\frac{1}{n}$ is the Euler constant :
 $\gamma=\lim_{n\to+\infty}\Big(\sum_{k=1}^{n}\frac{1}{k}-\log n\Big).$
@@
 $\int_{1}^{+\infty}b(t){f^{\prime}(t)}\ dt=i\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\,\ dt.$
-Thus, in this case the Ramanujan’s constant $C(f)$, previously defined in (2), is also
+Thus, in this case the Ramanujan constant $C(f)$, previously defined in (2), is also
 $C(f)=\frac{1}{2}f(1)+i\int_{0}^{+\infty}\frac{{f(1+it)-f(1-it)}}{e^{2\pi t}-1}\,\ dt.$ (4)
@@
 Let $f$ and $g$ two functions verifying the assumptions of theorem 1 and such that $f(n)=g(n)$ for every integer $n\geq 1$.
-Then Ramanujan’s constants $C(f)$ and $C(g)$, defined by (4), need not necessarily be the same.
+Then the Ramanujan constants $C(f)$ and $C(g)$, defined by (4), need not be the same.
 For example, if $f\equiv 0$ and $g:z\mapsto\sin(\pi z)$, then $C(f)=0$, but
 $C(g)=i\int_{0}^{+\infty}\frac{\sin\big(\pi(1+it)\big)-\sin\big(\pi(1-it)\big)}{e^{2\pi t}-1}\,\ dt=\int_{0}^{+\infty}\frac{e^{\pi t}-e^{-\pi t}}{e^{2\pi t}-1}\,\ dt\;\;=\;\int_{0}^{+\infty}e^{-\pi t}\ dt=\frac{1}{\pi}.$
-To avoid functions like $z\mapsto\sin(\pi z)$, we will consider a space of analytic functions such that Ramanujan’s constant, defined by (4), depends only on the values $f(n)$ for all integers $n\geq 1$}.
+To avoid functions like $z\mapsto\sin(\pi z)$, we will consider a space of analytic functions such that the Ramanujan constant, defined by (4), only depends on the values $f(n)$ for all integers $n\geq 1$}.
 2.3 The space $\mathcal{E}$
 Definition 1.
-Let $\mathcal{E}$} the space of analytic functions $f$ in the half plane $\{\Re(z)>0\}$ such that there is some constants $\emph{a}<\pi$ and $\emph{b}>0$ with :
+Let $\mathcal{E}$} the space of analytic functions $f$ in the half plane $\{\Re(z)>0\}$ such that there are some constants $\emph{a}<\pi$ and $\emph{b}>0$ with :
 for every ${\varepsilon}>0$, there is a constant $C_{\varepsilon}$ such that
@@
 The function $g\;:\;z\mapsto\sin(\pi z)$ is such that $g(n)=0$ for every integer $n\geq 1$}, but $g\notin\mathcal{E}$ since for $x\in\mathbb{R}$ and $t\in\mathbb{R}$
-$|g(x+it)|\sim\frac{1}{2}e^{\pi|t|}\text{ when }t\to\infty$
+$|g(x+it)|\sim\frac{1}{2}e^{\pi|t|}\text{ when }t\to\infty.$
 2.4 Ramanujan’s method of summation
@@
 2.5 Interpolation of partial sums
-To evaluate Ramanujan’s summation of a divergent series $\sum_{n\geq 1}f(n)$, a simpler way than the formula (6) can be obtained using an interpolation function of the partial sums of this series.
+To evaluate Ramanujan’s summation of a divergent series $\sum_{n\geq 1}f(n)$, a simpler formula than (6) can be obtained by means of an interpolation function of the partial sums of this series.
 In chapter VI of his Notebooks, Ramanujan uses, without any definition, such a function $\varphi$ which satisfies
@@
 Theorem 3.
-Let $f\in\mathcal{E}$. There is a unique function $\varphi_{f}\in\mathcal{E}$, such that
+Let $f\in\mathcal{E}$. There exists a unique function $\varphi_{f}\in\mathcal{E}$, such that
 […]
@@
 $\Re\zeta(\frac{1}{2}+it)$ $=$ $2\sum_{m=0}^{+\infty}x_{2m}Q_{2m}(t)$
-$\Im\zeta(\frac{1}{2}+it)$ $=$ $2\sum_{m=0}^{+\infty}x_{2m+1}Q_{2m+1}(t).$
+$\Im\zeta(\frac{1}{2}+it)$ $=$ $-2i\sum_{m=0}^{+\infty}x_{2m+1}Q_{2m+1}(t).$
 By (25) we have,